### v3

# Satisficing Regret Minimization in Bandits

[…]
Satisficing Regret Minimization in Bandits
[…]
Satisficing Regret Minimization in Bandits: Constant Rate and Light-Tailed Distribution
[…]
Motivated by the concept of satisficing in decision-making, we consider the problem of satisficing regret minimization in bandit optimization. In this setting, the learner aims at selecting satisficing arms (arms with mean reward exceeding a certain threshold value) as frequently as possible. The performance is measured by satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. We propose SELECT, a general algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence bound Testing, that attains constant expected satisficing regret for a wide variety of bandit optimization problems in the realizable case (i.e., a satisficing arm exists). Specifically, given a class of bandit optimization problems and a corresponding learning oracle with sub-linear expected (standard) regret upper bound, SELECT iteratively makes use of the oracle to identify a potential satisficing arm with low regret. Then, it collects data samples from this arm, and continuously compares the LCB of the identified arm’s mean reward against the threshold value to determine if it is a satisficing arm. As a complement, SELECT also enjoys the same (standard) regret guarantee as the oracle in the non-realizable case. To further ensure stability of the algorithm, we introduce SELECT-LITE that achieves a light-tailed satisficing regret distribution plus a constant expected satisficing regret in the realizable case and a sub-linear expected (standard) regret in the non-realizable case. Notably, SELECT-LITE can operate on learning oracles with heavy-tailed (standard) regret distribution. More importantly, our results reveal the surprising compatibility between constant expected satisficing regret and light-tailed satisficing regret distribution, which is in sharp contrast to the case of (standard) regret. Finally, we conduct numerical experiments to validate the performance of SELECT and SELECT-LITE on both synthetic datasets and a real-world dynamic pricing case study.
[…]
online learning, satisficing, regret, tail risk control
1 Introduction
Multi-armed bandit (MAB) is a classic framework for decision-making under uncertainty. In MAB, a learner is given a set of arms with initially unknown mean rewards. She repeatedly picks an arm to collect its noisy reward and her goal is to maximize the cumulative rewards. The performance is usually measured by the notion of regret, which is the difference between the maximum possible total mean rewards and the total mean rewards collected by the learner. With its elegant form, MAB has found numerous applications in recommendation systems, ads display, and beyond, and many (near-)optimal algorithms have been developed for various classes bandit optimization problems (Lattimore and Szepesvári 2020).
While making the optimal decision is desired in certain high-stakes situations, it is perhaps not surprising that what we often need is just a good enough or satisficing alternative (Lufkin 2021). Coined by Nobel laureate Herbert Simon (Simon 1956), the term “satisficing” refers to the decision-making strategy that settles for an acceptable solution rather than exhaustively hunting for the best possible.
The concept of satisficing finds its place in many real world decision-making under uncertainty settings (Caplin et al. 2011) as shown below.
  • Inventory Management: Studies show that product stockouts can result in customer disengagement and impair a firm’s profitability (Fitzsimons 2000, Jing and Lewis 2011). As such, retailers usually replenish their products periodically to improve service level (i.e., probability of no stockout for any product). Notably, due to various constraints (e.g., capacity and inventory cost), instead of blindly maximizing service level, retailers typically set a target service level, e.g., the ideal service level in retailing is $90\%\sim 95\%$ (Levy et al. 2023). Consequently, in face of the unknown demand distributions, one can cast the problem into a MAB setting with a different objective. In this setting, a retailer starts without complete knowledge of the demand distributions and proceeds by trying various stock level combinations (i.e., arms) of the products over time. Here, the objective is to hit the target service level as frequently as possible rather than maximizing the service level every time.
  • Product Design: In fashion product design, designers usually need to design a sequence of products for a group of customers with initially unknown preferences. Instead of demanding for the best design every time, designers only need to pick a design that can hopefully meet the expectation of a certain portion of the customers for most of the time (Mabey 2022, Whitenton 2014).
The above examples give rise to the study of satisficing bandits (Tamatsukuri and Takahashi 2019, Michel et al. 2023) where the objective is to hit a satisficing arm (i.e., an arm whose mean reward exceeds a certain threshold value) as frequently as possible. The corresponding performance measure is satisficing regret, which is the cumulative deficit of the chosen arm’s mean reward compared to the threshold. In particular, Michel et al. (2023) develops SAT-UCB, an algorithm for finite-armed satisficing bandits: In the realizable case (i.e., a satisficing arm exists), the algorithm attains a constant expected satisficing regret, i.e., independent of the length of the time horizon; Otherwise, in the non-realizable case (i.e., the threshold value is out of reach), it preserves the optimal logarithmic expected (standard) regret.
Despite these, existing algorithms usually impose that there is a clear separation between the threshold value and the largest mean reward among all non-satisficing arms, which is referred to as the satisficing gap. Moreover, the satisficing regret bounds usually scale inverse proportionally to the satisficing gap. On the other hand, it is evident that in many bandit optimization problem classes (e.g., Lipschitz bandits and concave bandits), the satisficing gap can simply be 0, making the bounds vacuous. Therefore, it is not immediately clear if SAT-UCB (Michel et al. 2023) or other prior results can go beyond finite-armed bandits and if constant satisficing regret is possible for general classes of bandit optimization problems.
1.1 Main Contributions
In this work, we propose SELECT, a novel algorithmic template for Satisficing REgret Minimization via SampLing and LowEr Confidence Bound Testing. More specifically:
  • We describe the design of SELECT in Section 3. For a given bandit optimization problem class and an associated learning oracle with sub-linear (but not necessarily optimal) expected standard regret guarantee, SELECT runs in rounds with the help of the oracle. At the beginning of a round, SELECT first applies the oracle for a number of time steps and randomly samples an arm from its trajectory as a candidate satisficing arm. Then, it conducts forced sampling by pulling this arm for a pre-specified number of times. Finally, it continues to pull this arm and starts to monitor the resulted lower confidence bound (LCB) of its mean reward. SELECT terminates the current round and starts the next once the LCB falls below the threshold value;
  • In Section 4, we establish that in the realizable case, i.e., when a satisficing arm exists, SELECT is able to achieve a constant expected satisficing regret. Notably, the expected satisficing regret bound has no dependence on the satisficing gap (see the forthcoming Remark 3 for a detailed discussion). Instead, it scales inverse proportionally to the exceeding gap, which measures the difference between the optimal reward and the threshold value, and is positive in general; Otherwise (i.e., when no satisficing arm exists), SELECT enjoys the same expected standard regret bound as the learning oracle.
  • In Section 5, we instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and demonstrate the corresponding satisficing and standard regret bounds. We also provide some discussions on the respective lower bounds;
  • While SELECT attains constant satisficing regret in expectation, its performance in individual realizations may not be stable, i.e., the probability of incurring a large satisficing regret is not sufficiently small. To address this, in Section 6, we consider the tail risk control in satisficing regret minimization. To confirm the necessity of new algorithms for light-tailed satisficing regret, we first prove that the satisficing regret distribution of SELECT can be heavy-tailed even if we use a learning oracle with light-tailed standard regret distribution. Then, we introduce SELECT-LITE, a modified version of SELECT to obtain a light-tailed satisficing regret distribution. Specifically, with an arbitrary learning oracle with sub-linear expected standard regret bound, SELECT-LITE enjoys the following desirable properties:
    • In the realizable case, SELECT-LITE attains a constant expected satsficing regret bound, and its satisficing regret distribution is light-tailed; and
    • In the non-realizable case, SELECT-LITE still maintains a sub-linear in $T$ standard regret bound.
    Notably, our results do not rely on any assumptions on the tail distribution of the learning oracle’s standard regret. In fact, SELECT-LITE is able to make use of a learning oracle with heavy-tailed standard regret to obtain a light-tailed satisficing regret distribution in the realizable case. Moreover, in Simchi-Levi et al. (2024), it is shown that any algorithm with a light-tailed standard regret distribution has to suffer a $\Omega(\text{poly}(T))$ instance-dependent expected standard regret, rather than a near-optimal $O(\log(T))$ instance-dependent expected standard regret. Because the constant expected satisficing regret bounds are typically instance-dependent, it becomes unclear if a light-tailed satisficing regret distribution will inevitably cause any dependence on $T$ in the expected satisficing regret. Nevertheless, our results show that achieving a light-tailed satisficing regret only worsens the expected satisficing regret by a constant. This thus reveals the compatibility between constant expected satisficing regret and light-tailed satisficing regret distribution. It also indicates a fundamental difference between the notions of standard regret and satisficing regret.
  • In Section 7, we conduct numerical experiments to demonstrate the performance of SELECT on synthetic instances of finite-armed bandits, concave bandits, and Lipschitz bandits. We use SAT-UCB, SAT-UCB+ (a heuristic with no regret guarantee proposed by Michel et al. 2023), and the respective learning oracle as the benchmarks. Our results reveal that in the realizable case, SELECT does achieve constant expected satisficing regret. Moreover, both its expected satisficing regrets (in the realizable cases) and standard regrets (in the non-realizable cases) either significantly outperform the benchmarks’ or are close to the best-performing ones’. We also test the performance of SELECT-LITE and compare it with SELECT. We show that SELECT-LITE has better performance in terms of tail distribution control compared to SELECT while still maintaining a constant expected satisficing regret comparable to SELECT. This indicates that SELECT-LITE can provide better protection against tail risks without incurring a significantly larger satisficing regret in expectation.
  • To test the performance of our algorithm in real-world instances, in Section 8, we also conduct numerical experiments using a case study of dynamic pricing calibrated from a real-world dataset. Using a real-world avocado sales dataset Kiggins (2018), we show that SELECT attains a constant expected satisficing regret in the realizable case. Furthermore, our results reveal that SELECT outperforms all other benchmarks both in the realizable and non-realizable cases.
  • A preliminary version of this paper (Feng et al. 2025) has appeared in the Proceedings of the 13th International Conference on Learning Representations (ICLR 2025). The current paper provides additional contributions in three directions.
    • First, we go beyond minimizing satisficing regret only in expectation and consider the tail risk control in satisficing regret minimization. Specifically, we adapt SELECT and establish SELECT-LITE, which is able to attain a light-tailed satisficing regret distribution (Theorem 5) while still maintaining a constant expected satisficing regret in the realizable case (Theorem 6) for a wide range of bandit problems;
    • Next, in Section 7.4, we test the empirical performance of SELECT-LITE using an instance of finite-armed bandit. The results show that SELECT-LITE is able to maintain a smaller tail probability in satisficing regret compared to SELECT while the expected satisficing regret of SELECT-LITE is comparable to that of SELECT;
    • Finally, in Section 8, we test our algorithm using a case study of dynamic pricing calibrated from a real-world avocado sales dataset. We observe that under this instance SELECT is able to significantly outperform all other benchmarks, demonstrating a strong empirical performance of our algorithm in practical bandit problems.
1.2 Related Works
Besides Michel et al. (2023), the most relevant works to ours are Bubeck et al. (2013), Garivier et al. (2019), Tamatsukuri and Takahashi (2019), all of which focus on finite-armed bandits. Bubeck et al. (2013) and Garivier et al. (2019) introduce constant regret algorithms with knowledge of the optimal reward. This can be viewed as a special case of satisficing where the threshold value is exactly the optimal mean reward. Hüyük and Tekin (2021) further extend the algorithm in Garivier et al. (2019) to multi-objective multi-armed bandit settings. Tamatsukuri and Takahashi (2019) also establishes constant regret for finite-armed satisficing bandits when there exists exactly one satisficing arm. As a follow-up paper of Michel et al. (2023), Hajiabolhassan and Ortner (2023) considers satisficing in finite-state MDP.
Among others, Russo and Van Roy (2022) is one of the first to introduce the notion of satisficing regret, but they focus on a time-discounted setting. Furthermore, in Russo and Van Roy (2022), the learner is assumed to know the difference between the values of the threshold and the optimal reward.
More broadly, the concept of satisficing has also been studied by Reverdy and Leonard (2014), Reverdy et al. (2017), Abernethy et al. (2016) in bandit learning. Both Reverdy and Leonard (2014) and Abernethy et al. (2016) focus on maximizing the number of times that the selected arm’s realized reward exceeds a threshold value. This is equivalent to a MAB problem that sets each arm’s mean reward as its probability of generating a value that exceeds the threshold. In view of this, the problems investigated in Reverdy and Leonard (2014) and Abernethy et al. (2016) are closer to conventional MABs. They are also different from Reverdy et al. (2017), Michel et al. (2023) and ours, which use the mean reward of the selected arms to compute the satisficing regret. We note that in Reverdy et al. (2017), a more general Bayesian setting is studied. In this setting, the satisficing regret also takes the learner’s belief that some arm is satisficing into account.
The notion of satisficing is also studied in the pure exploration setting (Locatelli et al. 2016, Mukherjee et al. 2017, Kano et al. 2019) as thresholding bandits. Their objective is to identify all arms with mean reward above a certain satisficing level with limited exploration budget. Unlike the satisficing regret we consider, the performance of their algorithm is measured by simple regret, i.e., the expected number of misidentification made by the final output. Note that any pure exploration algorithm (Audibert and Bubeck 2010) is able to identify the optimal arm or a satisficing arm with high probability after a number of steps. However, due to a small but positive error probability, subsequent exploitation of the identified arm will always incur linear satisficing regret, thus a simple explore-then-exploit approach is unable to obtain a constant satisficing regret bound. Similar observations have also been made in Michel et al. (2023).
Another line of research relevant to our work is tail risk control in bandits. The general goal of this line of research is to design bandit algorithms that achieve light-tailed regret distributions besides minimizing expected regret. Some early works studying the tail distribution of bandit algorithms include Audibert et al. (2009) and Salomon and Audibert (2011), who show that for UCB algorithms, the probability of regret greater than $c(\log(T))^{p}$ (where $p>1$ is fixed) decays only polynomially in $T$. Fan and Glynn (2024) are among the first to establish bandit algorithms with a light-tailed (standard) regret distribution. They focus on the finite-armed bandit setting and show that for informationally-optimized bandit algorithms such as the UCB algorithm and Thompson sampling, the probability of incurring a linear in $T$ regret is at least $\Omega(1/T)$. Some follow-up works include Simchi-Levi et al. (2023) and Simchi-Levi et al. (2024). They show that any finite-armed bandit algorithm with a $O(\log(T))$ instance-dependent expected regret bound has a heavy-tailed regret distribution. They also establish an optimal tradeoff between tail risk, worst-case expected regret bound and instance-dependent expected regret bound. In our paper, we consider tail risk control for satisficing regret minimization. We show that unlike tail risk control in standard regret minimization, where $O(\log(T))$ instance-dependent expected regret bound cannot coexist with light-tailed regret distribution, in satisficing regret minimization there exists algorithm that attains a constant expected satisficing regret and a light-tailed satisficing regret distribution.
2 Problem Formulation
In this section, we present the setup of our problem. Consider a class of bandit optimization problems $(\mathcal{A},\mathcal{R})$, where $\mathcal{A}$ is the arm set and $\mathcal{R}\subseteq\{r:\mathcal{A}\to[0,1]\}$ is the class of admissible reward functions. The learner is given a set of feasible arms $\mathcal{A}$ and a class of admissible reward functions $\mathcal{R}$, pulling arm $A\in\mathcal{A}$ in a time step brings a random reward with mean $r(A)$. The learner knows the underlying reward function $r(\cdot)$ belongs to the function class $\mathcal{R}$, but the exact $r(\cdot)$ is initially unknown to the learner.
In each time step $t=1,2,\ldots,T$, the learner pulls an arm $A_{t}\in\mathcal{A}$, and then observes a noisy reward $Y_{t}=r(A_{t})+\epsilon_{t}$, where $\epsilon_{t}$ is a conditionally 1-subgaussian random variable, i.e., $\mathbb{E}[\epsilon_{t}|A_{1},\epsilon_{1}\ldots,A_{t-1},\epsilon_{t-1},A_{t}]=0$ and $\mathbb{E}[e^{\lambda\epsilon_{t}}|A_{1},\epsilon_{1}\ldots,A_{t-1},\epsilon_{t-1},A_{t}]\leq e^{\lambda^{2}/2}$ holds for all $\lambda\in\mathbb{R}$.
Remark 1.
[…]
In this work, we define two notions of regret. First, with the presence of satisficing level $S$, we consider the satisficing regret introduced in Reverdy et al. (2017), Michel et al. (2023), which is the expected cumulative deficit of the chosen arm’s mean reward when compared to $S$, i.e.
$\texttt{Regret}_{S}=\sum_{t=1}^{T}\max\{S-r(A_{t}),0\}.$
Because it is possible that no arm achieves the satisficing level, we also follow the conventional bandit literature to define the (standard) regret, which measures the expected cumulative difference of between the chosen arm’s mean reward and the optimal mean reward. Specifically, let $A^{*}=\arg\max_{A\in\mathcal{A}}r(A),$ the standard regret is defined as
$\texttt{Regret}=T\cdot r(A^{*})-\sum_{t=1}^{T}r(A_{t}).$
We say that the setting is realizable if $r(A^{*})\geq S$; Otherwise, it is non-realizable. We recall that an arm $A$ is satisficing if $r(A)\geq S$; Otherwise, it is non-satisficing.
Remark 2.
In the definition of satisficing regret, we consider the cumulative deficit of mean reward compared to the satisficing threshold because ideally we would like the arm pull in every time step to be satisficing, thus we penalize every arm pull of non-satisficing arms. Even though in some time steps the decision maker may pull an arm with mean reward over the satisficing threshold, the excessive mean reward cannot compensate for the satisficing regret incurred in other time steps. With this definition of satisficing regret, the problem cannot be trivially solved by running a traditional bandit algorithm. Although any bandit algorithm will gradually converge to the optimal arm and achieve a satisficing total reward with long enough time horizon, as time horizon $T$ increase the number of arm pulls of non-satisficing arm will also increase, incurring a satisficing regret scaling in $T$. In the forthcoming Section 7 and Section 8, we also numerically show that directly running bandit algorithms can incur substantial satisficing regret.
Existing works (e.g., Michel et al. 2023) are mostly focused on satisficing regret minimization in finite-armed bandit settings, and usually derive expected satisficing regret bound that scales inversely proportional to the satisficing gap
$\Delta_{S}=\min\{S-r(A):r(A)<S\}.$
However, for bandit optimization problems with large and even infinitely many arms, $\Delta_{S}$ can approach 0 quickly. To address this issue, we define the notion of exceeding gap that captures the difference between the optimal reward and $S$, i.e.,
$\Delta_{S}^{*}=r(A^{*})-S.$
3 SELECT: Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing
In this section, we present our algorithmic template SELECT for general classes of bandit optimization problems. When the arm set $\mathcal{A}$ is not too large, one can collect enough data samples for every arm to estimate its mean reward and gradually abandon any non-satisficing arms (see, e.g., Garivier et al. 2019, Michel et al. 2023). In general, however, the arm set can be large and may even contain infinitely many arms (e.g., continuous arm set). Therefore, it becomes impossible to identify all non-satisficing arms. We thus take an alternative approach: The algorithm repeatedly locates a potentially satisficing arm with low (satisficing) regret. Then, it tests if this candidate arm is truly a satisficing arm or not. If not, it kicks off the next round of search.
To ensure fast detection of candidate arms with low regret for a given bandit optimization problem class $(\mathcal{A},\mathcal{R})$, SELECT makes use of a bandit algorithm for (standard) regret minimization on $(\mathcal{A},\mathcal{R})$. To this end, we assume access to a blackbox learning oracle that achieves sub-linear standard regret for the bandit optimization problem class $(\mathcal{A},\mathcal{R})$.
Condition 1.
For problem class $(\mathcal{A},\mathcal{R})$, there exists a sequence of learning algorithms ALG such that for any reward function $r\in\mathcal{R}$ and any time horizon $t\geq 2$, the expected regret of $\texttt{ALG}(t)$ is bounded by
$t\cdot r(A^{*})-\mathbb{E}\left[\sum_{s=1}^{t}r(A_{s})\right]\leq C_{1}t^{\alpha}\log(t)^{\beta}.$
where $C_{1}\geq 1,1/2\leq\alpha<1,\beta\geq 0$ are constants only dependent on the problem class and independent from the time horizon $t$.
We remark that Condition 1 only asks for sub-linearity in standard regret upper bounds rather than optimality. Therefore, it is an extremely mild condition and can be easily satisfied by a wide range of bandit optimization problem classes and the corresponding bandit learning oracles, e.g., finite-armed bandits, linear bandits, concave bandits, Lipschitz bandits, etc. We will demonstrate this in the forthcoming Section 5 and Section I of the Appendix.
With this, for a problem class $(\mathcal{A},\mathcal{R})$ and its sub-linear regret learning algorithm ALG, SELECT runs in rounds. In round $i$, it takes the following steps (see Fig. 1 for a illustration):
Figure 1: Illustration of a single round of SELECT
Step 1. Identify a Potentially Satisficing Arm: Let $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$, SELECT follows $\texttt{ALG}(t_{i})$ for the first $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil=\lceil 2^{i/\alpha}\rceil$ steps of round $i$ and records its arm selections. Then, it samples an arm $\hat{A}_{i}$ uniformly random from the trajectory of $\texttt{ALG}(t_{i})$ in this round. By virtue of ALG, we find an arm whose mean reward is at most $\tilde{O}(t^{\alpha-1}_{i})=\tilde{O}(\gamma_{i})$ below $r(A^{*})$ in expectation, and only $\tilde{O}(t^{\alpha}_{i})$ satisficing regret is incurred in this step in the realizable case [^1] We adopt the asymptotic notations $O(\cdot),\Omega(\cdot),$ and $\Theta(\cdot)$ as defined in (Cormen et al. 2009). When logarithmic factors are omitted, we write them as $\tilde{O}(\cdot),\tilde{\Omega}(\cdot),$ and $\tilde{\Theta}(\cdot).$. We note that in the realizable case, as $i$ increases, $\tilde{O}(\gamma_{i})$ will gradually become smaller than $\Delta_{S}^{*}$, meaning that the sampled arm $\hat{A}_{i}$ is more likely to be a satisficing arm and enjoys no satisficing regret;
Step 2. Forced Sampling on the Identified Arm: To validate if $\hat{A}_{i}$ is a satisficing arm or not, we need to collect enough data from pulling $\hat{A}_{i}$ before performing any statistical test (see the forthcoming Remark 3 for the reasons). However, when the arm set $\mathcal{A}$ is large, $\texttt{ALG}(t_{i})$ might only have pulled $\hat{A}_{i}$ for limited number of times. To circumvent this, we perform a forced sampling on $\hat{A}_{i}$ by pulling it for $T_{i}=\lceil\gamma^{-2}_{i}\rceil$ times;
Step 3. Lower Confidence Bound Tests: In this step, SELECT kicks off the LCB test as more data is collected from pulling $\hat{A}_{i}$. We use $k$ (initialized to 0) to denote the number of additional pieces of noisy reward data generated by pulling $\hat{A}_{i}$ in the current time step and $\hat{r}^{\text{tot}}_{i}$ to denote the running total reward collected from Step 2 and the current step so far. At the same time, it keeps comparing the LCB of $\hat{A}_{i}$’s mean reward, i.e., $\hat{r}^{\text{tot}}_{i}/(T_{i}+k)-\sqrt{4\log(T_{i}+k)/(T_{i}+k)}$, against the threshold value $S$ after acquiring every piece of new data. SELECT terminates this round and enters the next whenever the LCB is less than $S$.
The pseudo-code of SELECT is presented in Algorithm 1.
Remark 3 (Key Novelties of SELECT).
Our algorithm brings together several key ingredients, including sampling from oracle trajectories (Step 1), forced sampling to collect data (Step 2), and the LCB test (Step 3). In what follows, we further comment on the subtleties of each step.
Step 1. Prior works on satisficing (Garivier et al. 2019, Michel et al. 2023) mainly use uniform exploration to find candidates of satisficing arms. This approach could incur major satisficing regret when the arm set is large or even possibly infinite, as conducting uniform exploration on a large arm set may result in pulling a large number of non-satisficing arms. To bypass this challenge, in Step 1, we search for candidates using a bandit learning oracle with sub-linear standard regret. Hence, we are able to find candidates of satisficing arms without incurring large regret.
Step 2. One challenge from using the LCB test in Step 3 is that it is a more conservative design compared to prior works. In fact, if we directly enter Step 3 without Step 2, the LCB of $r(\hat{A_{i}})$ can easily fall below $S$ even if $r(\hat{A}_{i})\geq S$. This is because we may not have pulled $\hat{A}_{i}$ for sufficiently many times in Step 1 and the corresponding confidence interval might be large. As a result, the lower confidence bound of $r(\hat{A}_{i})$, which is the difference between empirical mean and confidence radius, can be significantly below $S$. This indicates, without the forced sampling in Step 2, major satisficing regret can be incurred due to frequent re-start of a new round. With the forced sampled data collected from Step 2, SELECT ensures that the width of the confidence interval is of order $\tilde{O}(T_{i}^{-1/2})=\tilde{O}(\gamma_{i})$ (i.e., same as $\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]$). Consequently, whenever $\gamma_{i}$ shrinks to well below $\Delta_{S}^{*}$, SELECT gradually becomes less likely to terminate a round (see the forthcoming Proposition 2);
Step 3. In Step 3, SELECT compares the LCB of $\hat{A_{i}}$’s mean reward against $S$ to determine if $A_{i}$ is a satisficing arm. This deviates from prior works that use UCB (Garivier et al. 2019) or empirical mean (Michel et al. 2023). However, it turns out that this is essential in achieving a constant expected satisficing regret that does not scale with $1/\Delta_{S}$, which can quickly explode in many cases. As we will see, once entering Step 3, SELECT can terminate a round within 1 additional time step in expectation when facing a non-satisficing arm (i.e., (6) in the proof of the forthcoming Proposition 1). In contrast, if it follows prior works to use UCB or empirical mean instead, the number of time steps required (and hence, the satisficing regret) will unavoidably scale with $1/\Delta_{S}$ (see e.g., Theorem 9 of Garivier et al. 2019 or Theorem 1 of Michel et al. 2023).
Algorithm 1 Satisficing Regret Minimization via Sampling and Lower Confidence Bound Testing (SELECT)
Input: time horizon $T$, satisficing level $S$
Set $\gamma_{i}\leftarrow 2^{-i(1-\alpha)/\alpha}$ for all $i=1,2,3\dots$
for round $i=1,2,\dots$ do
     Set $t_{i}\leftarrow\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil$
     Run $\texttt{ALG}(t_{i})$, let $\{\bar{A}_{s}\}_{s\in[t_{i}]}$ be the trajectory of arm pulls
     Sample $q$ uniformly at random from $\{1,2,\dots,t_{i}\}$ and set $\hat{A}_{i}\leftarrow\bar{A}_{q}$
     Set $T_{i}\leftarrow\lceil\gamma_{i}^{-2}\rceil$, $k\leftarrow 0$
     Let $t$ be the current time step, pull arm $\hat{A}_{i}$ for the next $T_{i}$ time steps, and set $\hat{r}_{i}^{\text{tot}}\leftarrow\sum_{s=t}^{t+T_{i}-1}Y_{s}$
     while $LCB(\hat{A}_{i})\geq S$ do
         Set $k\leftarrow k+1$
         Pull arm $\hat{A}_{i}$ and observe $Y_{t}$, where $t$ is the current time step
         Update $\hat{r}_{i}^{\text{tot}}\leftarrow\hat{r}_{i}^{\text{tot}}+Y_{t}$
         Set
$LCB(\hat{A}_{i})\leftarrow\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}$
     end while
end for
[…]
4 Regret Analysis
In this section, we analyze the regret bounds of SELECT. We begin by providing an upper bound of expected satisficing regret in the realizable case, i.e., $r(A^{*})\geq S$.
Theorem 1.
If $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{C_{1}^{\frac{1}{1-\alpha}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\frac{\alpha}{1-\alpha}}\cdot\textup{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right),\,C_{1}T^{\alpha}\cdot\textup{polylog}(T)\right\}.$ (1)
We remark that, from the first term on the RHS of (1), SELECT can achieve a constant (w.r.t. $T$) expected satisficing regret in the realizable case. Moreover, even when $T$ is relatively small and the constant expected satisficing regret guarantee is loose compared to the oracle’s regret bound, it is able to adapt to the oracle’s performance and achieve a sub-linear in $T$ bound on expected satisficing regret.
Proof Sketch..
A complete proof for Theorem 1 is provided in Section A. The proof relies on two critical results. In the first one, we show an upper bound on the expected satisficing regret of round $i$.
Proposition 1.
If $r(A^{*})\geq S$, then the expected satisficing regret incurred in round $i$ is bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
The proof of this proposition is provided in Section A.1 of the appendix. We also show that once SELECT runs for enough rounds so that $\gamma_{i}=\tilde{O}(\Delta_{S}^{*})$, it is unlikely to start a new round.
Proposition 2.
If $r(A^{*})\geq S$, then for every $i$ that satisfies
$\dfrac{32C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*}\,,$ (2)
the probability that round $i$ ends within the time horizon $T$ (conditioned on the event that round $i$ is started within the time horizon) is bounded by $1/4$.
The proof of this proposition is provided in Section A.2 of the appendix. Since each round of SELECT runs independently, Proposition 2 indicates that the total number of rounds for SELECT can be upper-bounded by a shifted geometric random variable with success rate $3/4$.
With Proposition 1 and Proposition 2, we are ready to finish the proof of Theorem 1. We denote $i_{0}$ as the minimum integer such that (2) holds. We prove that $2^{i_{0}(1-\alpha)/\alpha}=1/\gamma_{i_{0}}=C_{1}/\Delta_{S}^{*}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. We further have $i_{0}=\Theta(\log(C_{1}/\Delta_{S}^{*}))$. Therefore by Proposition 1 the expected satisficing regret incurred in the first $i_{0}$ rounds is bounded by
$\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\frac{\alpha}{1-\alpha}}\log(2/\gamma_{i})^{\beta}\leq C_{1}^{\frac{1}{1-\alpha}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\frac{\alpha}{1-\alpha}}\cdot\text{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right).$
Then we bound the expected satisficing regret incurred after round $i_{0}$. By Proposition 2 SELECT enters round $i_{0}+k$ with probability $4^{-k}$. Therefore the expected satisficing regret after round $i_{0}$ is
$\sum_{k=1}^{\infty}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i_{0}+k}^{-\frac{\alpha}{1-\alpha}}\log(2/\gamma_{i_{0}+k})^{\beta}\cdot 4^{-k}=\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i_{0}}^{-\frac{\alpha}{1-\alpha}}\sum_{k=1}^{\infty}2^{-k}\left(\dfrac{(1-\alpha)(i_{0}+k)}{\alpha}+1\right)^{\beta}\log(2)^{\beta}.$
We have shown that $\gamma_{i_{0}}^{-\alpha/(1-\alpha)}=(C_{1}/\Delta_{S}^{*})^{\alpha/(1-\alpha)}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. We further show that the summation above is polynomial in $i_{0}$, thus polynomial in $\log(C_{1}/\Delta_{S}^{*})$. Therefore the expected satisficing regret is upper bounded by $C_{1}^{1/(1-\alpha)}(1/\Delta_{S}^{*})^{\alpha/(1-\alpha)}\cdot\text{polylog}(C_{1}/\Delta_{S}^{*})$. Therefore the expected satisficing regret is upper bounded by the first term in the regret bound.
We further show that the expected satisficing regret is upper bounded by the second term in the regret bound. Let $i_{\max}$ be the total number of rounds for SELECT. We show that $t_{i_{\max}}=\gamma_{i_{\max}}^{-1/(1-\alpha)}=O(T)$, thus $i_{\max}=O(\log(T))$. By Proposition 1 the expected satisficing regret of each round is upper bounded by $C_{1}T^{\alpha}\cdot\text{polylog}(T)$. Since the total number of rounds is upper bounded by $O(\log(T))$, the expected satisficing regret of SELECT is also upper bounded by $C_{1}T^{\alpha}\cdot\text{polylog}(T)$. ∎
We also show that, in the non-realizable case, SELECT enjoys the same regret bound as the oracle in Condition 1 (The proof is provided in Section B of the appendix).
Theorem 2.
If $r(A^{*})<S$, the standard regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}]\leq C_{1}T^{\alpha}\cdot\textup{polylog}(T).$
Proof of Theorem 2 is provided in Appendix B. Theorem 2 implies that the expected standard regret of SELECT is sub-linear in $T$ in the non-realizable case. The expected standard regret bound of SELECT matches that of the learning oracle within logarithmic factors of $T$, indicating a similar performance guarantee of SELECT compared to the standard regret minimization algorithm in the corresponding class of bandit optimization problems.
With the above results, we establish that under a very general condition, SELECT does achieve constant expected satisficing regret in the realizable case without compromising the standard regret guarantee in the non-realizable case. Altogether, these empower SELECT the potential to become a general algorithm for decision-making under uncertainty.
5 Examples
In what follows, we instantiate SELECT to several popular classes of bandit optimization problems, including finite-armed bandits, concave bandits, and Lipschitz bandits. Along the way, we showcase how our results enable constant expected satisficing regret for bandits with large and even infinite arm set, which was not achieved by existing algorithms. We remark that the examples given here are non-exhaustive and similar results can be derived for other classes of bandit optimization problems such as linear bandits. In all three examples of this section, we assume the mean reward of each arm is in $[0,1]$.
1. Finite-Armed Bandits: Consider any instance of finite-armed bandit with $K$ arms, i.e., $\mathcal{A}=[K]$. In this case, both the UCB algorithm (see, e.g., Bubeck and Cesa-Bianchi 2012) and Thompson sampling (see, e.g., Agrawal and Goyal 2017) achieve a regret bound of $O(\sqrt{KT\log(T)})$. Combining them with Theorems 1 and 2, we have the following corollary. At the end of this section, we also provide some discussions on the lower bounds of the expcted satisficing regret.
Corollary 1.
By using either the UCB algorithm or Thompson sampling as ALG, if $r(A^{*})\geq S$, the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{K}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{K}{\Delta_{S}^{*}}\right),\,\sqrt{KT}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq\sqrt{KT}\cdot\textup{polylog}(T).$
Remark 4 (Comparison with Garivier et al. 2019, Michel et al. 2023).
Garivier et al. (2019), Michel et al. (2023) also provide algorithms for the expected satisficing regret in finite-armed bandit settings. While our expected satisficing regret bound is incomparable to Garivier et al. (2019), which attains a $O(K/\Delta_{S})$ expected satisficing regret, we provide major improvement over Michel et al. (2023), which achieves $O(K/\Delta_{S}+K/(\Delta_{S}^{*})^{2})$ expected satisficing regret. Compared to Michel et al. (2023), our regret bound remove the dependence on $\Delta_{S}$ and the additional $1/\Delta_{S}^{*}$ factor. We also point out that if we want to achieve the same expected satisficing regret as Garivier et al. (2019), we can simply change the LCB test in Step 3 to a UCB test.
2. Concave Bandits: Consider any instance of the concave bandit, i.e., $\mathcal{A}\subset\mathbb{R}^{d}$ is a bounded convex set, and $r(A)$ is a concave and $1$-Lipschitz continuous function defined on $\mathcal{A}$. Agarwal et al. (2011) gives an algorithm that enjoys $\text{poly}(d)\sqrt{T}\cdot\text{polylog}(T)$ regret. Together with Theorems 1 and 2, we have the following corollary.
Corollary 2.
By using the algorithm in Agarwal et al. (2011) as ALG, if $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{\textup{poly}(d)}{\Delta_{S}^{*}}\cdot\textup{polylog}\left(\dfrac{d}{\Delta_{S}^{*}}\right),\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq\textup{poly}(d)\sqrt{T}\cdot\textup{polylog}(T).$
3. Lipschitz Bandits: Consider any instance of the Lipschitz bandit in $d$ dimensions, i.e., $\mathcal{A}=[0,1]^{d}$, and $r(A)$ is an $L$-Lipschitz function (here Lipschitz functions are defined in the sense of $\infty$-norm). Bubeck et al. (2011) introduces a uniformly discretized UCB algorithm to achieve a $O(L^{d/(d+2)}T^{(d+1)/(d+2)}\sqrt{\log(T)})$ regret upper bound bound. Together with Theorem 1 and Theorem 2, we have the following corollary.
Corollary 3.
By using the UCB algorithm with uniform discretization in Bubeck et al. (2011) as ALG, if $r(A^{*})\geq S$, then the expected satisficing regret of SELECT is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{\dfrac{L^{d}}{(\Delta_{S}^{*}/2)^{d+1}}\cdot\textup{polylog}\left(\dfrac{L}{\Delta_{S}^{*}}\right),\,L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T)\right\};$
If $r(A^{*})<S$, the standard regret of SELECT is bounded by $\mathbb{E}[\texttt{Regret}]\leq L^{\frac{d}{d+2}}T^{\frac{d+1}{d+2}}\cdot\textup{polylog}(T).$
Remark 5.
Recall that $\Delta_{S}=\min\{S-r(A):r(A)<S\}$, one can easily verify that in the realizable case, $\Delta_{S}=0$ for both concave bandits and the Lipschitz bandits. As such, one cannot directly apply the results in Garivier et al. (2019), Michel et al. (2023) to acquire a constant expected satisficing regret. With the notion of exceeding gap $\Delta_{S}^{*}$ and SELECT, we establish constant expected satisficing regret bounds for these two classes of bandit optimization problems.
4. Lower Bounds: To complement our main results, we also present the expected satisficing regret lower bounds for finite-armed bandits and concave bandits. The first one is an expected satisficing regret lower bound for the finite-armed bandits.
Theorem 3 (Finite-Armed Bandits).
For every non-anticipatory learning algorithm $\pi$, every $\Delta>0$ and $T\geq 1/\Delta^{2}$, there exists an instance of two-armed bandit such that $r(A^{*})-S=\Delta$, and the expected satisficing regret incurred by $\pi$ is at least $\Omega(1/\Delta)$.
Theorem 3 establishes a $\Omega(1/\Delta_{S}^{*})$ lower bound on expected satisficing regret for finite-armed bandits. The lower bound in Theorem 3 matches the expected satisficing regret bound of SELECT in Corollary 1 up to a factor of $K$ and a logarithmic factor of $K/\Delta_{S}^{*}$. If we consider the number of arms $K$ as a constant, then the expected satisficing regret bound of SELECT for finite-armed bandits is near-optimal.
Remark 6.
In Michel et al. (2023), the authors adapt the results from Bubeck et al. (2013) (for MAB with known optimal mean reward) to establish a $\Omega(1/\Delta_{S})$ expected satisficing regret lower bound for finite-armed bandits. This is different than the one in Theorem3, which is $\Omega(1/\Delta)$ with $\Delta$ being the exceeding gap (i.e., $\Delta_{S}^{*}$). This difference originates from the lower bound instances. The lower bound instance in Bubeck et al. (2013) (when adapted to satisficing bandits) sets the threshold value to the optimal mean reward, i.e., $S=r(A^{*})\,,$ which makes both the expected satisficing regret and standard regret lower bounds to be $\Omega(1/\Delta_{S})$. In our lower bound proof, we allow $S$ to be smaller than $r(A^{*})$, which leads to a lower bound that depends on the difference between these two quantities.
Next, we provide an expected satisficing regret lower bound for bandits with concave rewards.
Theorem 4 (Bandits with Concave Rewards).
For every non-anticipatory learning algorithm $\pi$, every $\Delta>0$ and every $T\geq 1/\Delta^{2}$, there exists an instance of one-dimensional bandit with concave reward such that $r(A^{*})-S=\Delta$, and the expected satisficing regret incurred by $\pi$ is at least $\Omega(1/\Delta)$.
Theorem 4 establishes a $\Omega(1/\Delta_{S}^{*})$ lower bound on the expected satisficing regret for concave bandits. The lower bound established in Theorem 4 matches the expected satisficing regret upper bound of SELECT in Corollary 2 up to a polynomial factor of $d$ and a logarithmic factor of $d/\Delta_{S}^{*}$. If we consider the dimension $d$ as a constant, then the expected satisficing regret bound of SELECT for concave bandits is near-optimal.
6 Tail Risk Control in Satisficing Regret Minimization
While SELECT achieves a constant satisficing regret in expectation, it is unclear if it enjoys a light-tailed distribution on the satisficing regret (i.e., the realized satisficing regret is unlikely to be much larger than the expectation), a highly desirable property in maintaining the stability of an algorithm (Fan and Glynn 2024, Simchi-Levi et al. 2023). We say the satisficing regret of an algorithm is light-tailed if the probability of satisficing regret being greater than $x$ decays at an rate of $\exp(-x^{\zeta})$ for some $\zeta>0$ as $x$ increases. In other words, the satisficing regret distribution is light-tailed if there exists $\Lambda_{1},\Lambda_{2}$ such that for any $T$ and any $0\leq x\leq T$,
$\text{Pr}(\texttt{Regret}_{S}>x)\leq\Lambda_{1}\exp\left(-x^{\zeta}/\Lambda_{2}\right).$ (3)
Here $\Lambda_{1},\Lambda_{2}>0$ are parameters independent of $x$. We remark that $\Lambda_{1}$ and $\Lambda_{2}$ can depend on the length of the time horizon $T$, but we require that $\Lambda_{1}$ could scale at most polynomially with $T$, and $\Lambda_{2}$ could scale at most poly-logarithmically with $T$.
A light-tailed satisficing regret distribution is critical in many practical settings of satisficing regret minimization, particularly in applications sensitive towards tail risk such as healthcare and supply chain. In these applications, a light-tailed satisficing regret distribution reduces the possibility of suffering substantial losses for the decision maker, thus providing the decision maker with more protection against extreme scenarios. For instance, in the inventory management example discussed in Section 1, if an algorithm is able to achieve a light-tailed satisficing regret distribution, it indicates it is highly unlikely for the retailers to frequently fall short of the designated service level over time.
6.1 Key Challenges and Insufficiency of SELECT
Based on prior results (see, e.g., Fan and Glynn 2024, Simchi-Levi et al. 2023, Simchi-Levi et al. 2024), achieving a light-tailed satisficing regret distribution while still maintaining a constant expected satisficing regret can be extremely challenging due to the following reasons:
  1. 1.
    Tradeoff between Instance-Dependency and Tail Risk: Existing works show that in standard regret minimization for finite-armed bandits, a worse instance-dependent expected standard regret is inevitable for any algorithm with a light-tailed standard regret distribution. In particular, Simchi-Levi et al. (2024) prove that any algorithm with a light-tailed standard regret distribution has to suffer a $\Omega(\text{poly}(T))$ instance-dependent expected standard regret, rather than a near-optimal $O(\log(T))$ instance-dependent expected standard regret for the UCB algorithm and Thompson sampling. Since existing constant expected satisficing regret bounds (e.g., Garivier et al. 2019, Michel et al. 2023, and ours) are all instance-dependent, one may conjecture that in satisficing regret minimization, a light-tailed satisficing regret distribution could also inevitably come at a cost of worse (e.g., $\text{polylog}(T)$ or $\text{poly}(T)$) instance-dependent expected satisficing regret. Then, it becomes unclear whether a constant expected satisficing regret bound can coexist with a light-tailed satisficing regret distribution even in the finite-armed bandit setting.
  2. 2.
    Heave-Tailed Learning Oracle: To the best of our knowledge, existing works on tail risk control for standard regret focus mostly on finite-armed bandits and linear bandits, and it remains unclear whether there exists learning oracles with light-tailed regret distribution for more general classes of bandit optimization problems. Therefore, we may have to start with learning oracles with potentially heavy-tailed standard regret distribution. This makes our algorithm design even more challenging in achieving a light-tailed satisficing regret distribution for more general classes of bandit optimization problems.
Before proceeding, we first confirm the necessity of new algorithmic techniques in achieving a light-tailed satisficing regret distribution. We use the following proposition to show that the satisficing regret distribution of SELECT can indeed be heavy-tailed even if the learning oracle ALG used is already light-tailed.
Proposition 3.
For every finite-armed bandit learning oracle with $\tilde{O}(\sqrt{T})$ standard regret, if the oracle selects each arm at least once with probability at least $1/2$, then there exists an instance of two-armed bandit such that for all $T\in\mathbb{Z}_{+}$, the satisficing regret of SELECT satisfies
$\text{Pr}\left(\texttt{Regret}_{S}\geq\frac{T}{6}\right)\geq\frac{1}{16}\exp\left(-3\log(T)^{2}\right).$
The proof of Proposition 3 is provided in Section D of the appendix. Proposition 3 states that the satisficing regret distribution resulted from SELECT is heavy-tailed under mild conditions of the learning oracle. In fact, such conditions in Proposition 3 hold widely in most of the existing algorithms for finite-armed bandits, such as the UCB algorithm, Thompson sampling, and even for the algorithm in Simchi-Levi et al. (2024), whose standard regret distribution is known to be light-tailed. This suggests that simply applying SELECT does not guarantee a light-tailed satisficing regret distribution, even if the learning oracle adopted is light-tailed. Therefore, a new algorithm is needed if we want to ensure a light-tailed satisficing regret distribution.
6.2 SELECT-LITE: SELECT with Light-Tailed Satisficing Regret Distribution
In this section, we present SELECT-LITE, an augmented algorithmic template with a light-tailed satisficing regret distribution for more general classes of bandit optimization problems.
We note that there are two factors that prevent SELECT from achieving a light-tailed satisficing regret distribution. First, the length $t_{i}$ of Step 1 and the length $T_{i}$ of Step 2 increase exponentially as rounds proceed. Since larger $t_{i}$ and $T_{i}$ often result in larger tail risk, the rapid growth of $t_{i}$ and $T_{i}$ can lead to uncontrolled tail risk, eventually resulting in a heavy-tailed satisficing regret distribution. Second, in the LCB test in Step 3 of SELECT, we construct a sequence of confidence intervals for the candidate satisficing arm such that the $k$-th confidence interval contains the true mean reward with probability $1-\text{poly}(1/k)$. Although we have shown that these confidence intervals guarantee removal of a non-satisficing candidate arm in constant time steps in expectation, the distribution of the number of steps in the LCB test could also be heavy-tailed, potentially leading to a heavy-tailed satisficing regret distribution.
In observance of these, we make two modifications to SELECT to make the satisficing regret distribution light-tailed. First, in the modified algorithm, we set the length $t_{i}^{\prime}$ of Step 1 and the length $T_{i}^{\prime}$ of Step 2 to scale polynomially with the number of rounds $i$ instead of exponentially. Second, we use larger confidence intervals in the LCB test such that the $k$-th confidence interval contains the mean reward with probability at least $1-\exp(-\Omega(k^{\zeta}))$. With this configuration of confidence intervals, if the candidate satisficing arm is non-satisficing, then the probability that the LCB test lasts for more than $k$ time steps is at most $\exp(-\Omega(k^{\zeta}))$, thus we are able to guarantee that the number of steps for the LCB test in each round follows a light-tailed regret distribution.
We refer to this algorithm as SELECT-LITE, and the complete algorithm is provided in Algorithm 2. Here, $\Gamma(z)$ is the gamma function given by $\Gamma(z)=\int_{0}^{\infty}t^{z-1}\exp(-t)\text{d}t$.
Algorithm 2 Satisficing Regret Minimization with Light-Tailed Satisficing Regret Distribution (SELECT-LITE)
Input: time horizon $T$, satisficing level $S$, tail distribution parameter $\zeta$
Set $\gamma_{i}\leftarrow i^{-(1/\zeta-1)(1-\alpha)}$ for all $i=1,2,3\dots$
for round $i=1,2,\dots$ do
     Set $t^{\prime}_{i}\leftarrow\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil$
     Run $\texttt{ALG}(t^{\prime}_{i})$, let $\{\bar{A}_{s}\}_{s\in[t^{\prime}_{i}]}$ be the trajectory of arm pulls
     Sample $q$ uniformly at random from $\{1,2,\dots,t^{\prime}_{i}\}$ and set $\hat{A}_{i}\leftarrow\bar{A}_{q}$
     Set $T^{\prime}_{i}\leftarrow\lceil\gamma_{i}^{-2}\rceil$, $k\leftarrow 0$
     Let $t$ be the current time step, pull arm $\hat{A}_{i}$ for the next $T^{\prime}_{i}$ time steps, and set $\hat{r}_{i}^{\text{tot}}\leftarrow\sum_{s=t}^{t+T^{\prime}_{i}-1}Y_{s}$
     while $LCB(\hat{A}_{i})\geq S$ do
         Set $k\leftarrow k+1$
         Pull arm $\hat{A}_{i}$ and observe $Y_{t}$, where $t$ is the current time step
         Update $\hat{r}_{i}^{\text{tot}}\leftarrow\hat{r}_{i}^{\text{tot}}+Y_{t}$
         Set
$LCB(\hat{A}_{i})\leftarrow\dfrac{\hat{r}_{i}^{\text{tot}}}{T^{\prime}_{i}+k}-\sqrt{\frac{k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k}}$
     end while
end for
[…]
6.3 Tail Risk and Regret Analysis of SELECT-LITE
In this section, we provide a tail risk and regret analysis of SELECT-LITE. We show that SELECT-LITE has the following desirable properties: in the realizable case, SELECT-LITE enjoys both a light-tailed satisficing regret distribution and a constant expected satisficing regret bound at the same time; in the non-realizable case, SELECT-LITE still maintains a sub-linear standard regret bound.
We first use the following theorem to show that SELECT-LITE indeed has a light-tailed satisficing regret distribution in the realizable case.
Theorem 5.
If $r(A^{*})\geq S$, then for all $x>0$, the probability that satisficing regret of SELECT-LITE exceeds $x$ is bounded by
$\text{Pr}\left(\texttt{Regret}_{S}>x\right)\leq\exp\left(2i_{0}+6-\left(\frac{(x-x_{0})_{+}}{9+(1-\zeta)^{-1}(1-\alpha)^{-1}}\right)^{\zeta}\right)+\exp\left(2i_{0}+8-\frac{(x-x_{0})_{+}^{\zeta}}{6}\right),$
where $i_{0}=\Theta\left(\left(1/\Delta_{S}^{*}\cdot\log(1/\Delta_{S}^{*})^{\beta^{\prime}}\right)^{\frac{\zeta}{(1-\zeta)^{2}(1-\alpha)}}\right)$ and $x_{0}=6(i_{0}+\zeta(i_{0}+1)^{1/\zeta})$ are constants independent of $x$ and $T$.
Proof Sketch.
The complete proof of Theorem 5 is provided in Appendix E of the appendix. Similar to Proposition 2, we first prove that in SELECT-LITE, the probability of round $i$ terminating within the time horizon is at most $1/4$ for all large enough $i$.
Proposition 4.
If $r(A^{*})\geq S$, then for all $i$ that satisfies
$\dfrac{32C_{1}(1-\zeta)^{-\zeta/2}}{(1-\alpha)^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*},$ (4)
the probability that round $i$ ends within the time horizon $T$ (conditioned on the event that round $i$ is started within the time horizon) is bounded by $1/4$.
Proposition 4 implies that the total number of rounds for SELECT-LITE can be upper-bounded by a shifted geometric random variable with success rate $3/4$. Proof of Proposition 4 is provided in Section E.1 of the appendix.
The inequality in the theorem is trivial when $x<x_{0}$, thus we focus on the case where $x\geq x_{0}$. We denote $i_{0}$ as the smallest $i$ that satisfies (4). Since $x>x_{0}$, we get $x\geq 3\sum_{i=1}^{i_{0}}(t_{i}^{\prime}+T_{i}^{\prime})$. Then we divide the satisficing regret of SELECT-LITE into three parts: 1. the satisficing regret incurred in Step 1 and 2 by round $i_{0}$; 2. the satisficing regret incurred in Step 1 and 2 after round $i_{0}$; 3. the satisficing regret incurred in Step 3 of all rounds. Since $x\geq 3\sum_{i=1}^{i_{0}}(t_{i}^{\prime}+T_{i}^{\prime})$, the first part of satisficing regret is at most $x/3$. Therefore if the total satisficing regret exceeds $x$, then either the second or the third part of satisficing regret exceeds $x/3$.
To bound the tail distribution of the second part of satisficing regret, we first define $k_{0}$ as the maximum integer such that $\sum_{i=i_{0}+1}^{i_{0}+k_{0}}(t_{i}^{\prime}+T_{i}^{\prime})\leq x/3$. Then we verify that $k_{0}=\Omega(x^{\zeta})$. If the second part of satisficing regret exceeds $x/3$, then the total number of rounds should be at least $i_{0}+k_{0}+1$, and by Proposition 4, SELECT-LITE enters round $i_{0}+k_{0}+1$ with probability at most $4^{-k_{0}}=\exp(-\Omega(x^{\zeta}))$. Therefore the probability that the second part of satisficing regret exceeds $x/3$ is at most $\exp(-\Omega(x^{\zeta}))$.
To bound the tail distribution of the third part of satisficing regret, by the definition of confidence intervals we have that the probability that the number of steps in the LCB test exceeds $k$ is $O(\exp(-k^{\zeta}))$. Furthermore, by Proposition 4, the total number of rounds of SELECT-LITE is upper bounded by a shifted geometric random variable with success probability $3/4$. Then we use the following proposition to bound the tail distribution of the third part of satisficing regret.
Proposition 5.
Suppose $\{Q_{i}\}$ are independent nonnegative random variables and there exists $K_{1}>0$, $10<\zeta<$ 1 such that for all $x>0$ and all $i\in\mathbb{Z}^{+}$,
$\text{Pr}(Q_{i}\geq x)\leq K_{1}\exp(-x^{\zeta}).$
Let $N(\lambda)$ be a geometric distribution with success rate being $1-\lambda$. Then for all $x>0$
$\text{Pr}\left(\sum_{i=1}^{N+N(\lambda)}Q_{i}\geq x\right)\leq(1+K_{1})^{N}\exp\left(\frac{1+K_{1}}{1-\lambda}-\frac{x^{\zeta}}{2}\right).$
Proposition 5 states that the sum of a random number of independent light-tailed random variables is still light-tailed if the number of random variables follows a shifted geometric distribution. Proof of Proposition 5 is provided in Section E.2. By Proposition 5, the probability that the total number of time steps for Step 3 exceeds $x/3$ is at most $\exp(-\Omega(x^{\zeta}))$. Therefore the probability that the satisficing regret incurred in Step 3 exceeds $x/3$ is at most $\exp(-\Omega(x^{\zeta}))$. Therefore we conclude that the probability that the satisficing regret of SELECT-LITE exceeds $x$ is at most $\exp(-\Omega(x^{\zeta}))$ for large enough $x$. ∎
Remark 7 (Independence of $T$ for the Tail Probability Bound).
We would like to point out that the tail probability bound of SELECT-LITE does not depend on $T$. In other words, for any realizable instance, there exists a fixed light-tailed distribution that dominates the distribution of satisficing regret of SELECT-LITE with any time horizon $T$. This is in sharp contrast with the tail probability bounds provided in prior works such as Simchi-Levi et al. (2024), which deteriorate as $T$ increases although the standard regret distribution for every fixed $T$ is light-tailed.
We use the following theorem to show that SELECT-LITE enjoys a constant expected satisficing regret bound.
Theorem 6.
If $r(A^{*})\geq S$, then the expected satisficing regret of SELECT-LITE is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{C_{1}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right)^{\frac{\zeta+\alpha(1-\zeta)}{(1-\zeta^{2})(1-\alpha)}}\cdot\textup{polylog}\left(\dfrac{C_{1}}{\Delta_{S}^{*}}\right),\,C_{1}T^{\alpha(1-\zeta)+\zeta}\cdot\textup{polylog}(T)\right\}.$
Proof of Theorem 6 is similar to that of Theorem 1 and is provided in Appendix F of the appendix. Theorem 6 states that for all $10<\zeta<$ 1, SELECT-LITE attains a constant expected satisficing regret bound depending on $\zeta$. The power of $1/\Delta_{S}^{*}$ in the expected satisficing regret bound is monotonically increasing in $\zeta$, suggesting that stronger guarantees on tail distribution of satisficing regret leads to larger constant satisficing regret on its expectation. When $\zeta$ approaches zero, the power of $1/\Delta_{S}^{*}$ approaches $\alpha/(1-\alpha)$, which is the same as the one in the expected satisficing regret bound of SELECT. This implies that when $\zeta$ is small, SELECT-LITE can have a similar expected satisficing regret as SELECT.
Finally, we show that in the non-realizable case, the expected standard regret of SELECT-LITE is sub-linear in $T$.
Theorem 7.
If $r(A^{*})<S$, the expected standard regret of SELECT-LITE is bounded by
$\mathbb{E}[\texttt{Regret}]\leq C_{1}T^{\alpha(1-\zeta)+\zeta}\cdot\textup{polylog}(T).$
Proof of Theorem 7 is provided in Appendix G of the appendix. Theorem 7 states that for all $10<\zeta<$ 1, SELECT-LITE always attains a sub-linear in $T$ expected standard regret bound in the non-realizable case. When $\zeta$ approaches zero, the power of $T$ in the expected standard regret bound of SELECT-LITE approaches $\alpha$, which is the power of $T$ in the expected standard regret bound of SELECT and the learning oracle used in Step 1. In other words, when $\zeta$ is small, SELECT-LITE can have a similar expected standard regret as SELECT.
6.4 Novelty of the Results
In this section, we would like to highlight some major novelties of our algorithm SELECT-LITE.
New Approach towards Light-Tailed Regret Distribution: In SELECT-LITE, a light-tailed satisficing regret distribution is attained mainly from a novel multi-round structure of the algorithm, where we gradually increase the number of time steps of running the learning oracle to search for satisficing arms. In Proposition 4, we prove that the number of rounds of SELECT-LITE is upper bounded by a shifted geometric distribution, which implies that the probability of starting a new round decays exponentially as the rounds proceed. With a polynomial growth in the length of running the oracle and an exponential decay in the probability of starting a new round, we are able to prove that the total number of time steps of running the oracle and forced sampling. This design is novel in the literature of tail risk control in bandits. Our approach is also fundamentally different from existing ones in Fan and Glynn (2024) and Simchi-Levi et al. (2024), where a light-tailed regret distribution is obtained by using larger confidence intervals that hold with a probability of at least $1-\exp(-\Omega(T^{\zeta}))$.
Nuanced Difference between Satisficing Regret and Standard Regret: Simchi-Levi et al. (2024) show that for standard regret minimization in finite-armed bandits, any algorithm with a light-tailed standard regret distribution will inevitably incur a polynomial in $T$ instance-dependent expected standard regret. In other words, no algorithm can maintain both a light-tailed standard regret distribution and a near-optimal $O(\log(T))$ instance-dependent expected standard regret bound at the same time. Because the constant expected satisficing regret bounds are typically instance-dependent, it becomes unclear if a constant expected satisficing regret is compatible with a light-tailed satisficing regret distribution. In this paper, we show that it is indeed possible to break this impossible regime. Specifically, SELECT-LITE  is able to attain a light-tailed satisficing regret distribution and a constant instance-dependent expected satisficing regret bound in the realizable case. Furthermore, our results can be extended beyond finite-armed bandits and are applicable to any class of bandit optimization problems that has a learning oracle with a sub-linear expected standard regret bound. This reveals a fundamental difference between the notions of standard regret and satisficing regret.
No Additional Assumptions on the Learning Oracle: We emphasize that our results do not rely on any condition on the learning oracle other than a sub-linear expected standard regret (Condition 1). In particular, we do not require that the standard regret distribution of the learning oracle is light-tailed. Instead, SELECT-LITE is able to make use of learning oracles with a heavy-tailed standard regret distribution to attain a light-tailed satisficing regret distribution in the realizable case. For example, the satisficing regret distribution of SELECT-LITE is light-tailed when using the UCB algorithm or Thompson sampling as the learning oracle in Step 1, even though both algorithms are proven to have a heavy-tailed standard regret distribution (see Fan and Glynn 2024). The reason why the tail distribution of the learning oracle’s standard regret does not change the result is that with the help of Proposition 4, we are able to directly prove that the total number of time steps of running the learning oracle and forced sampling follows a light-tailed distribution. Since we have assumed that the mean reward of each arm belongs to $[0,1]$, the total satisficing regret incurred when running the learning oracle and forced sampling follows a light-tailed distribution.
The ability of making use of learning oracles with potentially heavy-tailed standard regret distributions is important for SELECT-LITE to generalize to a broader range of bandit optimization problem classes. To the best of our knowledge, algorithms with light-tailed standard regret distribution have been established for only a very limited range of bandit optimization problems including finite-armed bandits and linear bandits. By accepting algorithms with potentially heavy-tailed standard regret distribution as learning oracle, SELECT-LITE is able to extend beyond this scope to any class of bandit optimization problems that admits a learning oracle with sub-linear expected standard regret bound.
Remark 8 (Light-Tailed Standard Regret in the Non-Realizable Case).
We would like to clarify that although we have proven a light-tailed satisficing regret distribution for SELECT-LITE in the realizable case, SELECT-LITE may not be able to attain a light-tailed standard regret distribution in the non-realizable case. The main reason for this is that we make no assumption on the standard regret distribution of the learning oracle. Because in the non-realizable case, SELECT-LITE could run the learning oracle in $\Theta(T)$ time steps, SELECT-LITE could have a heavy-tailed standard regret distribution in the non-realizable case if the standard regret distribution of the learning oracle is heavy-tailed. In Appendix H of the appendix, we introduce an algorithm for finite-armed bandits that, besides achieving a constant expected satisficing regret, attains both a light-tailed satisficing regret distribution in the realizable case and a light-tailed standard regret distribution in the non-realizable case.
7 Numerical Experiments on Synthetic Data
In this section, we conduct numerical experiments to test the performance of SELECT on finite-armed bandits, concave bandits, and Lipschitz bandits. All experiments are run on a PC with 4-core CPU and 16GB of RAM.
7.1 Finite-Armed Bandits
Setup: In this case, we consider an instance of $4$ arms, and the expected rewards of all arms are $\{0.6,0.7,0.8,1\}$. We vary the length of the time horizon from $500$ to $5000$ with a stepsize of $500$. We conduct experiments for both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.93$; for the non-realizable case, we set the satisficing level $S=1.5$. In both cases, we compare SELECT against Thompson sampling in Agrawal and Goyal (2017) as well as SAT-UCB and SAT-UCB+ in Michel et al. (2023). For SELECT, we use Thompson sampling as the learning oracle. The experiment is repeated for 1000 times and we report the average satisficing regret in the realizable case and the average standard regret in the non-realizable case.
Results: The results of the realizable case are presented in Figure 2(a), and the results of the non-realizable case are presented in Figure 2(b). From Figure 2(a) one can see that the expected satisficing regret of SELECT, SAT-UCB and SAT-UCB+ barely increases when $T$ exceeds $2500$. This suggests that all three algorithms attain a constant expected satisficing regret in the realizable case. On the other hand, Thompson sampling fails to exhibit a constant expected satisficing regret in the realizable case. When comparing the expected satisficing regret of the four algorithms, one can see that SELECT incurs a smaller satisficing regret compared to Thompson sampling and SAT-UCB, and the expected satisficing regret of SELECT is comparable with that of SAT-UCB+.
From Figure 2(b) one can see that the expected standard regret of SELECT in the non-realizable case is comparable to that of Thompson sampling, and smaller than that of SAT-UCB and SAT-UCB+. The reason why SELECT outperforms SAT-UCB and SAT-UCB+ is that in the non-realizable case, both SAT-UCB and SAT-UCB+ largely follow the UCB algorithm, while SELECT largely follows Thompson sampling. Due to the stronger empirical performance of Thompson sampling compared to the UCB algorithm, SELECT is able to outperform both SAT-UCB and SAT-UCB+ in the non-realizable case.
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 2: Comparison of SELECT , Thompson sampling, SAT-UCB and SAT-UCB+
7.2 Concave Bandits
Setup: In this case, we consider an instance with arm set $[0,1]$ and concave reward function $r(x)=1-16(x-0.25)^{2}$. We vary the length of the time horizon from $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level $S=0.3$; for the non-realizable case, we set the satisficing level $S=1.5$. For both cases, we compare our algorithm SELECT with the convex bandit algorithm introduced in Agarwal et al. (2011). For SELECT, we use the algorithm in Agarwal et al. (2011) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm set uniformly with stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm set. The experiment is repeated for 1000 times, and we report the average results.
Results: The results of realizable case are provided in Figure 3(a), and the results of non-realizable case are provided in Figure 3(b). One can see from Figure 3(a) that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $500$, suggesting that SELECT attains a constant expected satisficing regret in the realizable case. On the other hand, all the other three algorithms fail to exhibit a constant expected satisficing regret in the realizable case. When comparing the expected satisficing regret of the four algorithms, we observe that SELECT incurs a smaller expected satisficing regret than all three other algorithms.
One can see from Figure 3(b) that in the non-realizable case, SELECT incurs less expected standard regret compared to the algorithm in Agarwal et al. (2011). Compared to SAT-UCB and SAT-UCB+, the expected standard regret of SELECT is larger, but the performance of SELECT is comparable with both SAT-UCB and SAT-UCB+.
In this experiment, the reason why SELECT outperforms the algorithm in Agarwal et al. (2011) can be explained as follows. In the instance we use in this experiment, both the set of optimal arms and the set of significantly suboptimal arms are large, and directly running the algorithm in Agarwal et al. (2011) over the entire time horizon tend to over-explore due to larger confidence intervals, thus would potentially incur significant regret. In contrast, SELECT starts with running the algorithm in Agarwal et al. (2011) over a small number of time steps $t_{i}$. This allows SELECT to have a much smaller confidence intervals in early stages of the algorithm, thus explores less and exploits more aggressively. Furthermore, when running forced sampling in Step 2, SELECT also exploits a reasonably good arm obtained from running the learning oracle, which means even more exploitation for SELECT. Therefore, although theoretically SELECT may underperform Agarwal et al. (2011) by a logarithmic factor of $T$ in the worst case, in the instance used in Figure 3(b), SELECT outperforms the algorithm in Agarwal et al. (2011). Similar phenomenon can also be observed in the numerical experiment for Lipschitz bandits in Section 7.3.
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 3: Comparison of  SELECT , Stochastic Convex, SAT-UCB and SAT-UCB+ for concave bandit
7.3 Lipschitz Bandits
Setup: In this case, we consider a two-dimensional Lipschitz bandit in domain $(x,y)\in[0,1]^{2}$ with Lipschitz rewards $f(x,y)=\min\{1,3e^{-100((x-0.5)^{2}+(y-0.7)^{2})}\}$. We vary the length of the time horizon from $500$ to $5000$ with a stepsize of $500$. We consider both the realizable case and the non-realizable case. For the realizable case, we set the satisficing level S = 0.5; for the non-realizable case, we set the satisficing level S = 1.5. For both cases, we compare our algorithm SELECT with the uniformly discretized UCB introduced in Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”). For SELECT, we use the Uniform UCB as the learning oracle. We again use SAT-UCB and SAT-UCB+ as heuristics by uniformly discretizing the arm set with stepsize $L^{-1/4}T^{-1/4}$, where $L$ is the Lipschitz constant of the reward function, and run the algorithms over the discretized arms. The experiment is repeated for 1000 times and we report the average results.
Results: The results of realizable case are provided in Figure 4(a), and the results of non-realizable case are provided in Figure 4(b). One can see from Figure 4(a) that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $500$, suggesting that SELECT attains a constant expected satisficing regret in the realizable case. On the other hand, none of the other three algorithms are able to attain a constant expected satisficing regret. When comparing the expected satisficing regret of the four algorithms, one can observe that SELECT incurs an expected satisficing regret smaller than all three other algorithms.
As for the non-realizable case, one can see from Figure 4(b) that SELECTachieves the best empirical performance among the four algorithms, while the empirical performance of SAT-UCB and SAT-UCB+ is similar to that of Uniform UCB in the non-realizable case.
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 4: Comparison of  SELECT , Uniform UCB, SAT-UCB and SAT-UCB+ for Lipschitz bandit
7.4 Tail Risk Control in Finite-Armed Bandits
We test the performance of SELECT-LITE in terms of both expected satisficing regret and tail distribution of satisficing regret using an instance of finite-armed bandit.
Setup: We consider an instance of 4 arms. The expected rewards of all arms are set as $\{0.2,0.4,0.6,0.8\}$. For the realizable case, we set the satisficing level $S=0.7$; for the non-realizable case, we set the satisficing level $S=1.5$. For SELECT and SELECT-LITE, we use Thompson sampling as the learning oracle. We vary $\zeta\in\{0.1,0.4\}$.
We first compare the expected satisficing regret of SELECT-LITE with that of SELECT and Thompson sampling. In this case, we vary the length of the time horizon from 1000 to 10000 with a stepsize of 1000. The experiment is repeated for 1000 times and we report the average satisficing regret. Then we compare the empirical tail distribution of satisficing regret between SELECT-LITE and SELECT under time horizon $T=10000$. The experiment is repeated for 10000 times.
Results: The results of realized satisficing regret distribution under $\zeta=0.1$ are presented in Figure 5(a) for SELECT and Figure 5(b) for SELECT-LITE , and the comparison on tail probability is presented in Figure 5(c). The results under $\zeta=0.4$ are presented in Figure 6. Here we report the empirical distribution of realized satisficing regret using histograms with a bin width of 10. From Figure 5(c) and 6(c), one can see that under both $\zeta=0.1$ and $\zeta=0.4$, SELECT-LITE exhibits a lower tail probability compared to SELECT. When comparing between between the case of $\zeta=0.1$ and the case of $\zeta=0.4$, SELECT-LITE with $\zeta=0.1$ exhibits a noticeably thinner tail, even though a larger $\zeta$ should theoretically yield the faster decay rate. This is because when $\zeta=0.1$, $t_{i}^{\prime}$ increases rapidly in the first few rounds, thus SELECT-LITE can quickly find the satisficing arm within a few rounds. Consequently, the empirical tail of SELECT-LITE when $\zeta=0.4$ is a summation of tails of many rounds, but the one when $\zeta=0.1$ is a summation of simply one or two rounds, which potentially makes the tail probability smaller.
The results of expected satisficing regret in the realizable case are presented in Figure 7(a). The results of expected standard regret in the non-realiable case are presented in Figure 7(b). From Figure 7(a) one can observe that SELECT-LITE  exhibits constant satisficing regret that is only slightly larger than that of SELECT. Furthermore, both when $\zeta=0.1$ and when $\zeta=0.4$, SELECT-LITE incur less satisficing regret compared to Thompson sampling. From Figure 7(b) one can see that SELECT-LITE  with different $\zeta$ incurs roughly the same standard regret as SELECT and is slightly better than Thompson sampling. This is because each SELECT-LITE algorithm shares a backbone with SELECT, and SELECT-LITE with $\zeta=0.1$ has a $O(T^{0.55})$ regret bound that is very close to that of SELECT. For SELECT-LITE with $\zeta=0.4$, although the theoretical standard regret is $O(T^{0.7})$, worse than the $O(T^{0.5})$ guarantee for Thompson sampling, it outperforms Thompson sampling empirically because its aggressive exploration yields a sizable constant-factor advantage over the time horizons studied. With these observaations, we conclude that by running SELECT-LITE instead of SELECT, the decision maker can have better protection against tail risks, both theoretically and empirically, without incurring a significantly larger satisficing regret in expectation.
(a) SELECT Regret Distribution
(b) SELECT-LITE Regret Distribution
(c) Tail Comparison
Figure 5: Realized satisficing regret distribution of SELECT  and SELECT-LITE  at $T=10000$ with $\zeta=0.1$
(a) SELECT Regret Distribution
(b) SELECT-LITE Regret Distribution
(c) Tail Comparison
Figure 6: Realized satisficing regret distribution of SELECT  and SELECT-LITE  at $T=10000$ with $\zeta=0.4$
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 7: Comparison of SELECT , SELECT-LITE  with different $\zeta$, and Thompson sampling
8 Numerical Experiments on Real Data: Satisficing in Dynamic Pricing
To further evaluate the empirical performance of our algorithm in more practical settings, in this section, we conduct numerical experiments on an instance calibrated from the avocado dataset Kiggins (2018).
8.1 Data Description, Pre-Processing, and Setup
We use the avocado dataset Kiggins (2018) in the New York region. The data records weekly average prices of organic avocado and the corresponding weekly total selling volume from January 4, 2015, to December 27, 2015, for a total of 52 weeks (with outliers deleted). The weekly average prices in these 52 data points range from \$1.65 to \$2.34, and the average price is \$2.03. The weekly selling volumes range from 9205 to 25675, and the average weekly selling volume is 17106.
We consider a dynamic pricing problem with linear demands to the prices, and test the performance of the algorithms on this model fitted by the avocado dataset. Assuming that the demand is equal to $D_{t}=g-hp_{t}+\epsilon_{t}$, where $\epsilon_{t}\sim N(0,\sigma^{2})$, given price $p_{t}$ at each time step $t$, we fit the linear regression model from the real data to be $\hat{g}=32724,\;\hat{h}=7678,\;\hat{\sigma}=4100.$ By normalizing the model to ensure unit noise variance, we obtain the coefficients to be
$\hat{g}_{n}=32724/4100,\;\hat{h}_{n}=7678/4100,\;\hat{\sigma}_{n}=1.$
We also extend the range of average price to be $p\in[0,4]$. The mean revenue received at each time $t$ is $p_{t}(\hat{g}_{n}-\hat{h}_{n}p_{t})$, and the mean revenue of each feasible price is visualized in Figure 8.
Figure 8: Expected revenue of the feasible prices $p\in[0,4]$
We set the length of the time horizon to be 500 to 8000 for the realizable case and 5000 for the non-realizable case with a stepsize of 500. For the realizable case, we set the satisficing level $S=8.0$; for the non-realizable case, we set the satisficing level $S=10.0$. For both cases, we compare our algorithm SELECT with Linear UCB introduced in Lattimore and Szepesvári (2020). For SELECT, we use Linear UCB in Lattimore and Szepesvári (2020) as the learning oracle. We also use SAT-UCB and SAT-UCB+ as heuristics by viewing the problem as Lipschitz bandits. That is, we first discretize the arm set uniformly with a stepsize $L^{-1/3}T^{-1/3}$, where $L$ is the Lipschitz constant of the reward function, and then run SAT-UCB and SAT-UCB+ over the discretized arm set. The experiment is repeated 1000 times and we report the averaged results.
8.2 Results
(a) Satisficing Regret in the Realizable Case
(b) Standard Regret in the Non-Realizable Case
Figure 9: Comparison of  SELECT , Linear UCB, SAT-UCB and SAT-UCB+ for linear bandit
The results of the realizable case are provided in Figure 9(a), and the results of the non-realizable case are provided in Figure 9(b). From Figure 9(a), one can observe that the expected satisficing regret of SELECT becomes stable after $T$ exceeds $3000$. This suggests that in the realizable case, SELECT is able to attain constant expected satisficing regret. On the other hand, none of the other three algorithms are able to achieve constant expected satisficing regret in the realizable case. When comparing the expected satisficing regret of the four algorithms, one can see that SELECT incurs a smaller expected satisficing regret than all the other three algorithms.
From Figure 9(b), one can see that in the non-realizable case, SELECT slightly outperforms Linear UCB and significantly outperforms both SAT-UCB and SAT-UCB+. The reason why SELECT and Linear UCB significantly outperforms SAT-UCB and SAT-UCB+ is that in the problem, SAT-UCB and SAT-UCB+ views the problem as a Lipschitz bandit, without making use of the linear structure of demand.
9 Conclusion
In this paper, we propose SELECT, a general algorithmic template, for satisficing regret minimization in bandit optimization. For a given class of bandit optimization problems and a sub-linear regret learning oracle, we show that SELECT attains constant expected satisficing regret in the realizable case and the same regret as the learning oracle in the non-realizable case. We instantiate SELECT to finite-armed bandits, concave bandits, and Lipschitz bandits, and also make a discussion on the corresponding lower bounds. We also consider tail risk control in satisficing regret minimization, where we introduce SELECT-LITE, which attains a light-tailed satisficing regret distribution while still maintaining a constant expected satisficing regret. Finally, we conduct numerical experiments to validate the performance of SELECT and SELECT-LITE on both synthetic datasets and a real-world dynamic pricing case study.
For future studies, we suggest further exploration in the following directions. First, our work mainly focuses on the stationary settings, and further research could explore satisficing in the nonstationary online learning settings (Wei and Luo 2021, Cheung et al. 2022). Also, the notion of satisficing regret indeed discourages exploration. Therefore, another direction is to study the role of satisficing regret in settings with limited adaptivity (Gao et al. 2019, Feng et al. 2023).
Appendix
Appendix A Proof of Theorem 1
Denote $\beta^{\prime}=\max\{\beta,1/2\}$ and $K_{\alpha,\beta}=\dfrac{32}{(1-\alpha)^{\beta}}$, and $i_{0}$ as
$i_{0}=\max\{i:C_{1}K_{\alpha,\beta}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}>\Delta_{S}^{*}\}.$
We first prove an upper bound on $i_{0}$. Denote $\delta_{0}\in(0,1)$ the smallest solution to the following equation
$\delta_{0}\log(2/\delta_{0})^{\beta^{\prime}}=\dfrac{\Delta_{S}^{*}}{C_{1}K_{\alpha,\beta}}.$
Then we have
$\dfrac{1}{\delta_{0}}=\dfrac{C_{1}K_{\alpha,\beta}\log(1/\delta_{0})^{\beta^{\prime}}}{\Delta_{S}^{*}}=\dfrac{(4\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}\log(2/\delta_{0}^{1/4\beta^{\prime}})^{\beta^{\prime}}}{\Delta_{S}^{*}}\leq\dfrac{(8\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}}{\Delta_{S}^{*}(\delta_{0})^{1/4}},$
or equivalently
$\dfrac{1}{\delta_{0}}\leq\left(\dfrac{(8\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}}{\Delta_{S}^{*}}\right)^{4/3},$
therefore
$\dfrac{1}{\delta_{0}}\leq\dfrac{(4/3)^{\beta^{\prime}}C_{1}K_{\alpha,\beta}\log({2(8\beta^{\prime})^{\beta^{\prime}}C_{1}K_{\alpha,\beta}}/{\Delta_{S}^{*}})^{\beta^{\prime}}}{\Delta_{S}^{*}}$
We write $D_{\alpha,\beta}=2(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta}$, $L_{\alpha,\beta}=(4/3)^{\beta^{\prime}}K_{\alpha,\beta}$, then we have
$\delta_{0}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta}C_{1}\log(D_{\alpha,\beta}C_{1}/\Delta_{S}^{*})^{\beta^{\prime}}}.$
By definition of $i_{0}$ we also have
$\gamma_{i_{0}}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta}C_{1}\log(D_{\alpha,\beta}C_{1}/\Delta_{S}^{*})^{\beta^{\prime}}},$
thus
$i_{0}\leq\dfrac{\alpha}{1-\alpha}\log_{2}\left(\dfrac{L_{\alpha,\beta}C_{1}\log(D_{\alpha,\beta}C_{1}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right).$
Using the result of Proposition 1 we have that the satisficing regret by round $i_{0}$ is bounded by
$\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\delta_{0})^{\beta}$
$=$ $\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}2^{i}\log(2/\delta_{0})^{\beta}$
$=$ $\dfrac{12C_{1}}{(1-\alpha)^{\beta}}2^{i_{0}}\log(2/\delta_{0})^{\beta}$
$\leq$ $\dfrac{16C_{1}}{(1-\alpha)^{\beta}}\left(\dfrac{L_{\alpha,\beta}C_{1}\log(D_{\alpha,\beta}C_{1}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)$
$=$ $\dfrac{16L_{\alpha,\beta}^{\alpha/(1-\alpha)}C_{1}^{1/(1-\alpha)}}{(1-\alpha)^{\beta}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)^{\alpha\beta^{\prime}/(1-\alpha)+1}.$
By the results of Proposition 2, for every $k\geq 1$, the probability that round $i_{0}+k$ starts within the time horizon is at most $4^{-k+1}$. Therefore the regret incurred after round $i_{0}$ is bounded by
$\sum_{k=1}^{\infty}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}2^{i_{0}+k}\log(2/\gamma_{i})^{\beta}\cdot 4^{-k+1}$
$\leq$ $\dfrac{12C_{1}}{(1-\alpha)^{\beta}}2^{i_{0}}\log(2/\delta_{0})^{\beta}\cdot\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$
$\leq$ $\dfrac{8L_{\alpha,\beta}^{\alpha/(1-\alpha)}C_{1}^{1/(1-\alpha)}M_{\beta}}{(1-\alpha)^{\beta}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)^{\alpha\beta^{\prime}/(1-\alpha)+1},$
where $M_{\beta}=\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$ is a constant only dependent on $\beta$, and the total satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\dfrac{8L_{\alpha,\beta}^{\alpha/(1-\alpha)}C_{1}^{1/(1-\alpha)}(M_{\beta}+2)}{(1-\alpha)^{\beta}}\left(\dfrac{1}{\Delta_{S}^{*}}\right)^{\alpha/(1-\alpha)}\log\left(\dfrac{D_{\alpha,\beta}C_{1}}{\Delta_{S}^{*}}\right)^{\alpha\beta^{\prime}/(1-\alpha)+1}.$
Furthermore, since the length of the time horizon is $T$ time steps, we have $t_{i}\leq T$, thus the maximum number of rounds is bounded by $i_{\max}=\lceil\alpha\log_{2}(T)\rceil$. By the results of Proposition 1 we have that the satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}$
$\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}2^{i+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{24C_{1}}{(1-\alpha)^{\beta}}T^{\alpha}\log(2T)^{\beta}.$
Combining the two parts of results we have completed the proof of Theorem 1.
A.1 Proof of Proposition 1
Recall that $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil\,,$ the satisficing regret incurred when running $\texttt{ALG}(t_{i})$ is bounded by
$\mathbb{E}\left[\sum_{s=1}^{t_{i}}\max\{S-r(A_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq$ $C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}$
$\leq$ $\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since $\hat{A}_{i}$ is selected randomly according to the trajectory of $\texttt{ALG}(t_{i})$ in round $i$, we further have
$\mathbb{E}[\max\{S-r(\hat{A}_{i}),0\}]\leq\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]=$ $\frac{\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]}{t_{i}}$
$\leq$ $\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ for $T_{i}=\lceil\gamma_{i}^{-2}\rceil$ times is bounded by
$\left\lceil\dfrac{1}{\gamma_{i}^{2}}\right\rceil\cdot\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-1}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
If $r(\hat{A}_{i})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{A}_{i}$ for $k$ times (in addition to the first $T_{i}$ times) occurs with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\dfrac{1}{(T_{i}+k)^{2}}\leq\dfrac{1}{(k+8)^{2}}.$ (5)
Therefore expected number of arm pulls of $\hat{A}_{i}$ after first pulling it $T_{i}$ times is bounded by
$\mathbb{E}\left[\text{Number of times }\hat{A}_{i}\text{ is pulled in round }i\text{ after the first }T_{i}\text{ times}\right]$
$=$ $\sum_{s=1}^{\infty}\text{Pr}(k\geq s)=\sum_{s=1}^{\infty}\dfrac{1}{(s+8)^{2}}\leq 1.$ (6)
Summing up the three parts of satisficing regret we have that the satisficing regret of round $i$ is bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
A.2 Proof of Proposition 2
We denote $A^{*}=\arg\max_{A\in\mathcal{A}}r(A)$.
By Condition 1 and recall from the proof of Proposition 1, we have
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]=$ $\dfrac{1}{t_{i}}\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq\dfrac{1}{t_{i}}C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}=C_{1}t_{i}^{\alpha-1}\log(t_{i})^{\beta}$
$\leq$ $\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta},$
where the last inequality holds because $t_{i}=\lceil 1/\gamma^{1/(1-\alpha)}\rceil\leq 2/\gamma^{1/(1-\alpha)}$.
According to the condition of this proposition, we have $i$ satisfies the condition
$\dfrac{32C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*},$
by the Markov’s inequality we have
$\text{Pr}(r(A^{*})-r(\hat{A}_{i})\geq\Delta_{S}^{*}/2)\leq\text{Pr}\left(r(A^{*})-r(\hat{A}_{i})\geq\dfrac{16C_{1}}{(1-\alpha)^{\beta}}\log(2/\gamma_{i})^{\beta}\right)\leq\dfrac{1}{8}.$ (7)
Now, if $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$, which happens with probability at least 7/8 according to the above inequality, then $r(\hat{A}_{i})\geq S+\Delta_{S}^{*}/2$ holds. On the other hand, if $i$ satisfies the condition stated in the proposition, then $\gamma_{i}\log(1/\gamma_{i})^{1/2}\leq\Delta_{S}^{*}/16$.
Note that $f(y)=\log(y)/y$ is monotonically decreasing when $y\geq 3$ and $32\cdot\log(2)/2\geq\log(3)/$ 3, for all $k\geq 0$,
$\sqrt{\dfrac{{4\log(T_{i}+k)}}{T_{i}+k}}\leq 4\sqrt{\dfrac{\log(T_{i})}{T_{i}}}\leq\dfrac{\Delta_{S}^{*}}{4}.$
Suppose round $i$ is terminated for some $k$, then
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}<S=r(A^{*})-\Delta_{S}^{*}\leq r(\hat{A}_{i})-\dfrac{\Delta_{S}^{*}}{2}\leq r(\hat{A}_{i})-2\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}},$
or equivalently
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}.$
By the Hoeffding’s inequality we have
$\text{Pr}\left(\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}\right)\leq\dfrac{1}{(T_{i}+k)^{2}}\leq\dfrac{1}{(k+4)^{2}}.$
Therefore conditioned on $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$ the probability that round $i$ is terminated by the end of the horizon is bounded by
$\sum_{k=1}^{\infty}\dfrac{1}{(k+8)^{2}}\leq\dfrac{1}{8}.$ (8)
Combining (7) and (8) we conclude that the probability that round $i$ terminates by the end of the time horizon is bounded by $1/4$.
Appendix B Proof of Theorem 2
We first prove that the regret incurred in round $i$ is at most
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
The regret incurred when running $\texttt{ALG}(t_{i})$ is bounded by
$\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq C_{1}t_{i}^{\alpha}\log(t_{i})^{\beta}\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
We further have
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ is bounded by
$\left\lceil\dfrac{1}{\gamma_{i}^{2}}\right\rceil\cdot\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}=\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-1}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since $r(\hat{A}_{i})\leq r(A^{*})<S$, for every $k$, round $i$ is not terminated with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}-\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}+k}--\sqrt{\dfrac{4\log(T_{i}+k)}{T_{i}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\dfrac{1}{(T_{i}+k)^{2}}\leq\dfrac{1}{(k+8)^{2}}.$
Therefore expected number of arm pulls of $\hat{A}_{i}$ after pulling $T_{i}$ times is bounded by
$\sum_{k=1}^{\infty}\dfrac{1}{(k+8)^{2}}\leq 1.$
Summing up the three parts of regret we have that the satisficing regret of round $i$ is bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since the length of the time horizon is $T$ time steps, we have $t_{i}\leq T$, thus the maximum number of rounds is bounded by $i_{\max}=\lceil\alpha\log_{2}(T)\rceil$. Therefore we have that the regret is bounded by
$\mathbb{E}[\texttt{Regret}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}$
$\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}2^{i+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{24C_{1}}{(1-\alpha)^{\beta}}T^{\alpha}\log(2T)^{\beta}.$
Therefore the regret is bounded by $C_{1}T^{\alpha}\cdot\text{polylog}(T)$.
Appendix C Proofs for Lower Bounds
C.1 Proof of Theorem 3
Denote $T_{0}=\lceil 1/\Delta^{2}\rceil$. We consider the following instances of two-armed bandit:
[…]
We further set the satisficing level $S=0$, and assume that noise $\epsilon_{t}$ follows i.i.d. standard normal distribution. Then for any $t\in[T_{0}]$,
$\texttt{KL}\left(\text{Pr}_{t}^{(1)}||\text{Pr}_{t}^{(2)}\right)=(t-1)\,\texttt{KL}(N(\Delta,1)||N(-\Delta,1))=\dfrac{t\Delta^{2}}{2}\leq\dfrac{\Delta^{2}}{2}\cdot\dfrac{1}{\Delta^{2}}=\dfrac{1}{2}.$
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$. Then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári 2020), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}\geq$ $\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(A_{t}=2)+\text{Pr}_{t}^{(2)}(A_{t}=1)$
$\geq$ $T_{0}\Delta\cdot\exp\left(-\texttt{KL}\left(\text{Pr}_{t}^{(1)}||\text{Pr}_{t}^{(2)}\right)\right)\geq\Delta\cdot\dfrac{1}{\Delta^{2}}\cdot\exp\left(-\dfrac{1}{2}\right)$
$=$ $\Omega\left(\dfrac{1}{\Delta}\right).$
Therefore any non-anticipatory learning algorithm will incur at least $\Omega(1/\Delta)$ satisficing regret under one of the two instances.
C.2 Proof of Theorem 4
Assume without loss of generality that $\Delta\leq 1$. Denote $T_{0}=\lceil 1/16\Delta^{2}\rceil$. We consider the following instances of two-armed bandit: $\mathcal{A}=[-3,3]$, and reward function is defined as follows:
[…]
We further set satisficing level $S=1-\Delta$, and assume that the noise $\epsilon_{t}$ follows i.i.d. standard normal distribution.
It is easy to verify that $|r^{(1)}(x)-r^{(2)}(x)|$ obtains its maximum at $-2\sqrt{2\Delta}$ and $2\sqrt{2\Delta}$, and
$|r^{(1)}(\sqrt{2\Delta})-r^{(2)}(\sqrt{2\Delta})|=2\sqrt{2\Delta}\cdot 2\sqrt{2\Delta}=8\Delta.$
Therefore for any $t\in[T_{0}]$,
$\texttt{KL}\left(\text{Pr}_{t}^{(1)}||\text{Pr}^{(2)}_{t}\right)\leq(t-1)\cdot\dfrac{(8\Delta)^{2}}{2}\leq\dfrac{1}{16\Delta^{2}}\cdot 32\Delta^{2}=2.$
We denote $\texttt{Regret}_{S}^{(i)}$ as the satisficing regret under instance $i$, then by the Bretagnolle-Huber inequality (see, e.g., Theorem 14.2 of Lattimore and Szepesvári 2020), we have
$\texttt{Regret}_{S}^{(1)}+\texttt{Regret}_{S}^{(2)}$ $\geq\Delta\sum_{t=1}^{T_{0}}\text{Pr}_{t}^{(1)}(A_{t}\geq 0)+\text{Pr}_{t}^{(2)}(A_{t}\leq 0)$
$\geq T_{0}\Delta\cdot\exp\left(-\texttt{KL}\left(\text{Pr}_{t}^{(1)}||\text{Pr}_{t}^{(2)}\right)\right)$
$\geq T_{0}\Delta\cdot\exp(-2)=\dfrac{\exp(-2)}{16\Delta^{2}}\cdot\Delta=\Omega(1/\Delta).$
Therefore any non-anticipatory learning algorithm will incur at least $\Omega(1/\Delta)$ in one of the two instances.
Appendix D Proof of Proposition 3
We consider a simple instance of two-armed bandit: $r(1)=1$ and $r(2)=0$ with $S=0.5$. Recall that the length of the time horizon is $T$ time steps, we have $t_{i}\leq T$, thus the maximum number of rounds is bounded by $i_{\max}=\lceil\alpha\log_{2}(T)\rceil$.
Consider that case that the total number of rounds is $l\leq i_{\max}$, and recall that the probability that arm 2 is pulled at least once in each round $i\in[l]$ is at least $1/2$. Define event :=$\mathcal{E}:=\{\text{Arm 2 is pulled in Step 2 for every round }i\in[l]\}$, then it happens with probability at least
[…]
$\geq\prod_{i=1}^{l}\frac{1}{2}\cdot\frac{1}{t_{i}}\geq\prod_{i=1}^{i_{\max}}\frac{1}{2}\cdot\frac{1}{t_{i}}\geq 2^{-(1+\log_{2}(T))}\cdot\prod_{i=1}^{1+\alpha\log_{2}(T)}2^{-(1+i/\alpha)}$
$=\frac{1}{4T^{2\alpha}}\cdot 2^{-(1+\alpha\log_{2}(T))^{2}/\alpha}\geq\frac{\exp(-\alpha\log(T)\log_{2}(T))}{16T^{2\alpha+2}}\geq\frac{1}{16}\exp(-3\log(T)^{2}).$
Condition on $\mathcal{E}$, if the total number of time steps for Step 3 is at least $T/6$, then the satisficing regret is at least $T/3$. Otherwise, the total number of time steps for Step 1 and 2 is at least $32T/$ 3. When $\alpha=1/2$, $t_{i}=\lceil 2^{2i}\rceil=T_{i}$ for every round before the final one. In the final round, the number of steps in Step 1 is also more than that in Step 2. Therefore the total number of time steps for Step 2 is at least $T/3$, which incurs a satisficing regret at least $T/3$. Therefore, we can bound the probability of total satisficing regret exceeding $T/6$ as
$\text{Pr}\left(\texttt{Regret}_{S}\geq\frac{T}{6}\right)\geq\text{Pr}\left(\texttt{Regret}_{S}\geq\frac{T}{6}\,\big{|}\,\mathcal{E}\right)\text{Pr}(\mathcal{E})\geq\text{Pr}(\mathcal{E})\geq\frac{1}{16}\exp\left(-3\log(T)^{2}\right).$
Recall that $\Lambda_{1}$ is at most polynomial in $T$ and $\Lambda_{2}$ is at most poly-logarithmically in $T$. We can state the heavy tail as the negation of the definition of light-tailed satisficing regret distribution in Equation 3 as
$\frac{\Lambda_{1}\exp(-T^{\zeta}/\Lambda_{2})}{\exp(-3\log(T)^{2})/16}=16\Lambda_{1}\exp(3\log(T)^{2}-T^{\zeta}/\Gamma_{2})$
$\leq$ $\exp(\text{polylog}(T))\cdot\exp\left(\text{polylog}(T)-\frac{T^{\zeta}}{\text{polylog}(T)}\right)<1,$
where the final inequality is because $\Omega(T^{\zeta})$ increases much faster than $\text{polylog}(T)$. Namely, for any $\Lambda_{1}\leq\text{poly}(T)$ and any $\Lambda_{2}\leq\text{polylog}(T)$, there exists $x=T/6$ such that
$\text{Pr}\left(\texttt{Regret}_{S}\geq\frac{T}{6}\right)\geq\frac{1}{16}\exp\left(-3\log(T)^{2}\right)>\Lambda_{1}\exp(-x^{\zeta}/\Lambda_{2}).$
Appendix E Proof of Theorem 5
The inequality in the theorem is trivial when $x<x_{0}$, thus for the rest of the proof we assume $x\geq x_{0}$. We first bound the probability that cumulative satisficing regret incurred in the Step 1 and 2 after round $i_{0}$ exceeds $x/3$. Define the satisficing regret incurred in the Step 1 and 2 in each round $i$ as $R_{i}$, then former probability can be divided into
$\text{Pr}\left(\sum_{i=i_{0}+1}^{\infty}R_{i}>\frac{x}{3}\right)$ $=\sum_{k=1}^{\infty}\text{Pr}\left(\sum_{i=i_{0}+1}^{\infty}R_{i}>\frac{x}{3}\bigg{|}\text{End at round }i_{0}+k\right)\cdot\text{Pr}(\text{Endat round }i_{0}+k)$
$\leq\sum_{k=1}^{\infty}\text{Pr}\left(\sum_{i=i_{0}+1}^{i_{0}+k}R_{i}>\frac{x}{3}\right)\cdot 4^{1-k}$
Denote $\beta^{\prime}=\max\{\beta,1/2\}$ and $K_{\alpha,\beta,\zeta}=\dfrac{32C_{1}(1-\zeta)^{-\zeta/2}}{(1-\alpha)^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}$, and $i_{0}$ as
$i_{0}=\max\{i:K_{\alpha,\beta,\zeta}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}>\Delta_{S}^{*}\}.$
We first prove an upper bound on $i_{0}$. Denote $\delta_{0}\in(0,1)$ the smallest solution to the following equation
$\delta_{0}^{1-\zeta}\log(2/\delta_{0})^{\beta^{\prime}}=\dfrac{\Delta_{S}^{*}}{K_{\alpha,\beta,\zeta}}.$
Then we have
$\dfrac{1}{\delta_{0}^{1-\zeta}}=\dfrac{K_{\alpha,\beta,\zeta}\log(1/\delta_{0})^{\beta^{\prime}}}{\Delta_{S}^{*}}=\dfrac{(4\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}\log(2/\delta_{0}^{1/4\beta^{\prime}})^{\beta^{\prime}}}{\Delta_{S}^{*}}\leq\dfrac{(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}{\Delta_{S}^{*}(\delta_{0})^{1/4}},$
or equivalently
$\dfrac{1}{\delta_{0}^{1-\zeta}}\leq\left(\dfrac{(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}{\Delta_{S}^{*}}\right)^{4/3},$
therefore
$\dfrac{1}{\delta_{0}^{1-\zeta}}\leq\dfrac{(4(1-\zeta)/3)^{\beta^{\prime}}K_{\alpha,\beta,\zeta}\log({2(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}/{\Delta_{S}^{*}})^{\beta^{\prime}}}{\Delta_{S}^{*}}$
We write $D_{\alpha,\beta,\zeta}=2(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}$, $L_{\alpha,\beta,\zeta}=(4(1-\zeta)/3)^{\beta^{\prime}}K_{\alpha,\beta,\zeta}$, then we have
$\delta_{0}^{1-\zeta}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}.$
By definition of $i_{0}$ we also have
$\gamma_{i_{0}}^{1-\zeta}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}},$
thus
$i_{0}\leq\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\zeta}{(1-\zeta)^{2}(1-\alpha)}}.$
We set
$k_{0}=\left\lfloor\left(\frac{x/3}{2+\zeta+(2(1/\zeta-1)(1-\alpha))^{-1}}\right)^{\zeta}-i_{0}-1\right\rfloor,$
then for any round $k\leq k_{0}$, the cumulative satisficing regret in Step 1 and 2 can be bounded as
$\sum_{i=i_{0}+1}^{i_{0}+k}R_{i}$ $\leq\sum_{i=i_{0}+1}^{i_{0}+k}(t_{i}^{\prime}+T_{i}^{\prime})\leq 2k+\sum_{i=i_{0}+1}^{i_{0}+k}(i^{1/\zeta-1}+i^{2(1/\zeta-1)(1-\alpha)})$
$<2k+\int_{i_{0}+2}^{i_{0}+k+1}(y^{1/\zeta-1}+y^{2(1/\zeta-1)(1-\alpha)})\mathrm{d}y$
$=2k+\left(\zeta y^{1/\zeta}+\frac{1}{2(1/\zeta-1)(1-\alpha)+1}y^{2(1/\zeta-1)(1-\alpha)+1}\right)\bigg{|}_{i_{0}+2}^{i_{0}+k+1}$
$\leq 2k+\left(\left(\zeta+\frac{1}{2(1/\zeta-1)(1-\alpha)}\right)y^{1/\zeta}\right)\bigg{|}_{i_{0}+2}^{i_{0}+k+1}$
$\leq\left(2+\zeta+\frac{1}{2(1/\zeta-1)(1-\alpha)}\right)(i_{0}+k+1)^{1/\zeta}\leq\frac{x}{3},$
where the final inequality is based on $k<k_{0}$ and the definition of $k_{0}$. Therefore if the cumulative satisficing regret in Step 1 and 2 exceeds $x/3$, then the total number of rounds should be at least $i_{0}+k_{0}+1$. Furthermore, by Proposition 4 we have that the probability that the total number of rounds exceeds $i_{0}+k_{0}$ is at most $4^{1-k_{0}}$. Therefore the probability of exceeding $x/3$ in Step 1 and 2 can be further bounded as
$\text{Pr}\left(\sum_{i=i_{0}+1}^{\infty}R_{i}>\frac{x}{3}\right)$ $4^{1-k_{0}}\leq{4^{i_{0}+3}}\cdot\exp\left(-\left(\frac{x/3}{2+\zeta+(2(1/\zeta-1)(1-\alpha))^{-1}}\right)^{\zeta}\right)$
$\leq\exp\left(6+2i_{0}-\left(\frac{x}{9+(1-\zeta)^{-1}(1-\alpha)^{-1}}\right)^{\zeta}\right).$
Then we bound the probability that cumulative satisficing regret incurred in the Step 3 of all rounds exceeds $x/3$. For each round $i$, the probability that round $i$ is not terminated after pulling non-satisficing arm $\hat{A}_{i}$ for $k_{i}$ times (in addition to the first $T_{i}^{\prime}$ times) occurs with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\dfrac{k_{i}^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k_{i}}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\dfrac{k_{i}^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k_{i}}}>r(\hat{A}_{i})\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k_{i}^{\zeta}\right).$ (9)
The probability that satisficing regret incurred in the Step 3 in round $i$ exceeds $x/3$ can be bounded by
$\text{Pr}\left(\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/3\right)\leq\text{Pr}\left(k_{i}>x/3\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-(x/3)^{\zeta}\right).$
Note that $\zeta/\Gamma(\zeta^{-1})\leq 3/2$, $\exp\left(3^{\zeta}\right)\leq e^{3}$. Using Proposition 5, the probability of total regret in LCB test exceeds $x/3$ can be bounded by
$\text{Pr}\left(\sum_{i=1}^{\infty}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/3\right)=\text{Pr}\left(\sum_{i=1}^{i_{0}+N(1/4)}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/3\right)$
$\leq$ $\left(1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(3^{\zeta}\right)\right)^{i_{0}}\exp\left(\frac{4}{3}\left(1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(3^{\zeta}\right)\right)-\frac{(x/3)^{\zeta}}{2}\right)\leq\exp\left(2i_{0}+8-\frac{x^{\zeta}}{6}\right)$
As we assume $x$ large enough with
$\sum_{i=1}^{i_{0}}(t_{i}^{\prime}+T_{i}^{\prime})<\sum_{i=1}^{i_{0}}2t_{i}^{\prime}<2i_{0}+\sum_{i=1}^{i_{0}}2i^{1/\zeta-1}$
$<$ $2i_{0}+2\int_{0}^{i_{0}}y^{1/\zeta-1}\mathrm{d}y=2i_{0}+2\zeta y^{1/\zeta}\bigg{|}_{1}^{i_{0}+1}<2i_{0}+2\zeta(i_{0}+1)^{1/\zeta}<x/3,$
namely $x>3\sum_{i=1}^{i_{0}}(t_{i}^{\prime}+T_{i}^{\prime})$, the satisficing regret incurred before round $i_{0}$ in Step 1 and 2 are always bounded by $x/3$. Recall $i_{0}$ is a constant without $x$ or $T$, we can bound the tail of total regret as
$\text{Pr}\left(\texttt{Regret}_{S}>x\right)\leq$ $\text{Pr}\left(\sum_{i=1}^{i_{0}}R_{i}>\frac{x}{3}\right)+\text{Pr}\left(\sum_{i=i_{0}+1}^{\infty}R_{i}>\frac{x}{3}\right)+\text{Pr}\left(\sum_{i=1}^{\infty}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/3\right)$
$\leq$ $\exp\left(6+2i_{0}-\left(\frac{x}{9+(1-\zeta)^{-1}(1-\alpha)^{-1}}\right)^{\zeta}\right)+\exp\left(2i_{0}+8-\frac{(x/3)^{\zeta}}{2}\right).$
Finally, since $(x-x_{0})_{+}\leq x$, by replacing $x$ with $(x-x_{0})_{+}$ we obtain the probability bound in the theorem.
E.1 Proof of Proposition 4
By Condition 1 we have
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]=\dfrac{1}{t_{i}^{\prime}}\mathbb{E}\left[t_{i}^{\prime}r(A^{*})-\sum_{s=1}^{t_{i}^{\prime}}r(A_{s})\right]\leq\dfrac{1}{t_{i}^{\prime}}C_{1}t_{i}^{\prime\alpha}\log(t_{i}^{\prime})^{\beta}\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta},$
where the last inequality holds because $t_{i}^{\prime}=\lceil 1/\gamma^{1/(1-\alpha)}\rceil\leq 2/\gamma^{1/(1-\alpha)}$.
According to the condition of this proposition, we have $i$ satisfies the condition
$\dfrac{32C_{1}(1-\zeta)^{-\zeta/2}}{(1-\alpha)^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*}.$
Note that for any $10<\zeta<$ 1, $(1-\zeta)^{-\zeta/2}>1$ and $\gamma_{i}^{1-\zeta}>\gamma_{i}$. Therefore $i$ satisfies
$\dfrac{32C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*},$
by the Markov’s inequality we have
$\text{Pr}(r(A^{*})-r(\hat{A}_{i})\geq\Delta_{S}^{*}/2)\leq\text{Pr}\left(r(A^{*})-r(\hat{A}_{i})\geq\dfrac{16C_{1}}{(1-\alpha)^{\beta}}\log(2/\gamma_{i})^{\beta}\right)\leq\dfrac{1}{8}.$ (10)
Now, if $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$, which happens with probability at least 7/8 according to the above inequality, then $r(\hat{A}_{i})\geq S+\Delta_{S}^{*}/2$ holds. On the other hand, if $i$ satisfies the condition stated in the proposition, then
$(1-\zeta)^{-\zeta/2}\cdot\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(1/\gamma_{i})^{1/2}\leq\Delta_{S}^{*}/16.$
Note that for $f(y)=y^{\zeta}/(T_{i}^{\prime}+y)$, it reaches its maximum when
$\frac{\mathrm{d}}{\mathrm{d}y}f(y)=\frac{\zeta y^{\zeta-1}T_{i}^{\prime}+y^{\zeta}(\zeta-1)}{(T_{i}^{\prime}+y)^{2}}=0.$
Therefore the maximum of $f(y)$ is reached at $\zeta T_{i}^{\prime}/(1-\zeta)$. For simplicity, we set $D=\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)$. The confidence radius for any $k$ is bounded by
$\sqrt{\frac{k^{\zeta}+D}{T_{i}^{\prime}+k}}\leq\sqrt{\frac{\max\{D,2\}k^{\zeta}}{T_{i}^{\prime}+k}}\leq\sqrt{\frac{\max\{D,2\}\left(\frac{\zeta T_{i}^{\prime}}{1-\zeta}\right)^{\zeta}}{T_{i}^{\prime}+\left(\frac{\zeta T_{i}^{\prime}}{1-\zeta}\right)}}$
$\leq$ $\sqrt{\frac{\max\{D,2\}(1-\zeta)^{-\zeta}\cdot\zeta^{\zeta}T_{i}^{\prime\zeta}}{T_{i}^{\prime}}}\leq\max\left\{\sqrt{D},2\right\}(1-\zeta)^{-\zeta/2}\gamma_{i}^{1-\zeta}\leq\frac{\Delta_{S}^{*}}{16}.$
Suppose round $i$ is terminated for some $k$, then
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}^{\prime}+k}-\sqrt{\frac{k^{\zeta}+D}{T_{i}+k}}<S=r(A^{*})-\Delta_{S}^{*}\leq r(\hat{A}_{i})-\dfrac{\Delta_{S}^{*}}{2}\leq r(\hat{A}_{i})-2\sqrt{\frac{k^{\zeta}+D}{T_{i}+k}},$
or equivalently
$\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}^{\prime}+k}\leq r(\hat{A}_{i})-\sqrt{\frac{k^{\zeta}+D}{T_{i}+k}}.$
By the Hoeffding’s inequality we have
$\text{Pr}\left(\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}^{\prime}+k}\leq r(\hat{A}_{i})-\sqrt{\frac{k^{\zeta}+D}{T_{i}^{\prime}+k}}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore conditioned on $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$ the probability that round $i$ is terminated by the end of the horizon is bounded by
$\sum_{k=1}^{\infty}\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$ (11)
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}.$
Combining (10) and (11) we conclude that the probability that round $i$ terminates by the end of the time horizon is bounded by $1/4$.
E.2 Proof of Proposition 5
We define $Y_{i}=Q_{i}^{\zeta}$ for all $i$, then for all $i\in\mathbb{Z}^{+}$,
$\text{Pr}(Y_{i}\geq x)=\text{Pr}(Q_{i}\geq x^{1/\zeta})\leq K_{1}\exp(-x).$
For all $i\in\mathbb{Z}^{+}$, we have
$\text{Pr}(\exp(Y_{i}/2)\geq x)=\text{Pr}(Y_{i}\geq 2\log(x))\leq K_{1}x^{-2}.$
Since $Q_{i}\geq 0$, we have $\exp(Y_{i}/2)\geq 1$, therefore
$\mathbb{E}[\exp(Y_{i}/2)]=\int_{0}^{\infty}\text{Pr}(\exp(Y_{i}/2)\geq x)\text{d}x\leq 1+\int_{1}^{\infty}K_{1}x^{-2}\text{d}x=1+K_{1}.$
We define $W=\sum_{i=1}^{N+N(\lambda)}Y_{i}$, and define $Z_{i}$ are independent Bernoulli random variables such that $\text{Pr}(Z_{i})=\lambda^{i}$. Then we have
$\mathbb{E}[\exp(W/2)]=(\mathbb{E}[\exp(Y_{1}/2)])^{N}\cdot\prod_{i=1}^{\infty}\mathbb{E}[\exp(Y_{1}Z_{i}/2)].$
By definition of $Z_{i}$ we have
$\mathbb{E}[\exp(Y_{i}Z_{i}/2)]=(1-\lambda^{i})+\lambda^{i}\mathbb{E}[\exp(Y_{1}/2)]\leq 1+(1+K_{1})\lambda^{i}.$
Therefore
$\mathbb{E}[\exp(W/2)]$ $\leq(1+K_{1})^{N}\cdot\prod_{i=1}^{\infty}(1+(1+K_{1})\lambda^{i})\leq(1+K_{1})^{N}\cdot\prod_{i=1}^{\infty}\exp((1+K_{1})\lambda^{i})$
$=(1+K_{1})^{N}\exp\left((1+K_{1})\sum_{i=1}^{\infty}\lambda^{i}\right)=(1+K_{1})^{N}\exp\left(\frac{1+K_{1}}{1-\lambda}\right).$
By Markov’s inequality we have
$\text{Pr}(W\geq x)=\text{Pr}(\exp(W/2)\geq\exp(x/2))\leq(1+K_{1})^{N}\exp((1+K_{1})/(1-\lambda)-x/2).$
Since $\zeta<1$, we have
$W=\sum_{i=1}^{N+N(\lambda)}Y_{i}=\sum_{i=1}^{N+N(\lambda)}Q_{i}^{\zeta}\geq\left(\sum_{i=1}^{N+N(\lambda)}Q_{i}\right)^{\zeta}.$
Therefore for any $x>0$,
$\text{Pr}\left(\sum_{i=1}^{N+N(\lambda)}Q_{i}\geq x\right)\leq\text{Pr}(W\geq x^{\zeta})\leq(1+K_{1})^{N}\exp\left(\frac{1+K_{1}}{1-\lambda}-\frac{x^{\zeta}}{2}\right).$
Appendix F Proof of Theorem 6
We first use the following proposition to bound the expected satisficing regret of each round.
Proposition 6.
If $r(A^{*})\geq S$, then the satisficing regret incurred in round $i$ is bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Proof of Proposition 6 is provided in Section F.1. Using the result of Proposition 6 we have that the satisficing regret by round $i_{0}$ is bounded by
$\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\delta_{0})^{\beta}=\sum_{i=1}^{i_{0}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}i^{\alpha(1-\zeta)/\zeta}\log(2/\delta_{0})^{\beta}$
$\leq$ $\dfrac{12C_{1}}{(1-\alpha)^{\beta}}i_{0}^{1+\alpha(1-\zeta)/\zeta}\log(2i_{0}^{(1/\zeta-1)(1-\alpha)})^{\beta}$
$\leq$ $\frac{12C_{1}}{(1-\alpha)^{\beta}(1-\zeta)}\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\zeta+\alpha(1-\zeta)}{(1-\zeta^{2})(1-\alpha)}}\log\left(\frac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta}$
By the results of Proposition 4, for every $k\geq 1$, the probability that round $i_{0}+k$ starts within the time horizon is at most $4^{-k+1}$. Therefore the regret incurred after round $i_{0}$ is bounded by
$\sum_{k=1}^{\infty}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}(i_{0}+k)^{\alpha(1-\zeta)/\zeta}\log(2/\gamma_{i})^{\beta}\cdot 4^{-k+1}$
$\leq$ $\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\left(\sum_{k=1}^{i_{0}}(2i_{0})^{\alpha(1-\zeta)/\zeta}\log(2/\gamma_{i})^{\beta}\cdot 4^{-k+1}+\sum_{k=i_{0}+1}^{\infty}(2k)^{\alpha(1-\zeta)/\zeta}\log(2/\gamma_{i})^{\beta}\cdot 4^{-k+1}\right)$
$\leq$ $\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\left(i_{0}\cdot(2i_{0})^{\alpha(1-\zeta)/\zeta}\log(2i_{0}^{(1/\zeta-1)(1-\alpha)})^{\beta}+\sum_{k=1}^{\infty}(2k)^{\alpha(1-\zeta)/\zeta}\log(2k^{(1/\zeta-1)(1-\alpha)})^{\beta}\cdot 4^{-k+1}\right)$
$\leq$ $\dfrac{6C_{1}\cdot 2^{\alpha(1-\zeta)/\zeta}}{(1-\alpha)^{\beta}}\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\zeta+\alpha(1-\zeta)}{(1-\zeta^{2})(1-\alpha)}}\log\left(\frac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta}+M,$
where $M=\sum_{k=1}^{\infty}(2k)^{\alpha(1-\zeta)/\zeta}\log(2k^{\alpha(1-\zeta)\zeta})^{\beta}\cdot 4^{-k+1}$ is a constant only dependent on $\alpha,\beta,\zeta$, and the total satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq M+$ $\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\left(2^{\alpha(1-\zeta)/\zeta}+\frac{2}{1-\zeta}\right)\left(\dfrac{L_{\alpha,\beta,\zeta}^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\zeta+\alpha(1-\zeta)}{(1-\zeta^{2})(1-\alpha)}}$
$\cdot\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\frac{\beta^{\prime}(\zeta+\alpha(1-\zeta))}{(1-\zeta^{2})(1-\alpha)}}\log\left(\frac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta}.$
Furthermore, since the length of the time horizon is $T$ time steps, we have
$T\geq\sum_{i}^{i_{\max}}t_{i}^{\prime}\geq\sum_{i}^{i_{\max}}i^{1/\zeta-1}\geq\int_{0}^{i_{\max}}x^{1/\zeta-1}\mathrm{d}x=\zeta i_{\max}^{1/\zeta},$
thus the maximum number of rounds is bounded by $i_{\max}=\lceil(T/\zeta)^{\zeta}\rceil$. By the results of Proposition 6 we have that the satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}\leq\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}i^{\alpha(1-\zeta)/\zeta}\log(2T)^{\beta}$
$\leq$ $\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\log(2T)^{\beta}\int_{1}^{i_{\max}+1}x^{\alpha(1-\zeta)/\zeta}\mathrm{d}x\leq\dfrac{6C_{1}}{(1-\alpha)^{\beta}(\alpha(1-\zeta)/\zeta+1)}(i_{\max}+1)^{\alpha(1-\zeta)/\zeta+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{6C_{1}\zeta^{-\alpha(1-\zeta)-\zeta}\cdot 2^{\alpha(\zeta^{-1}-1)+1}}{(1-\alpha)^{\beta}(\alpha(\zeta^{-1}-1)+1)}\,T^{\alpha(1-\zeta)+\zeta}\log(2T)^{\beta}$
F.1 Proof of Proposition 6
Following the proof pattern of Proposition 1, as the expression of $t_{i}^{\prime}$ in $\gamma_{i}$ remains $t_{i}^{\prime}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil,$ the satisficing regret incurred when running $\texttt{ALG}(t_{i}^{\prime})$ is still bounded by
$\mathbb{E}\left[\sum_{s=1}^{t_{i}^{\prime}}\max\{S-r(A_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}^{\prime}}r(A_{s})\right]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since $\hat{A}_{i}$ is selected randomly according to the trajectory of $\texttt{ALG}(t_{i}^{\prime})$ in round $i$, we still have
$\mathbb{E}[\max\{S-r(\hat{A}_{i}),0\}]\leq\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ for $T_{i}^{\prime}=\lceil\gamma_{i}^{-2}\rceil$ times is still bounded by
$\left\lceil\dfrac{1}{\gamma_{i}^{2}}\right\rceil\cdot\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-1}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
If $r(\hat{A}_{i})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{A}_{i}$ for $k$ times (in addition to the first $T_{i}^{\prime}$ times) occurs with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\frac{k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\frac{k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\exp\left(-k^{\zeta}+\log\left(\frac{\zeta}{8\Gamma(\zeta^{-1})}\right)\right)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore expected number of arm pulls of $\hat{A}_{i}$ after first pulling it $T_{i}^{\prime}$ times is bounded by
$\mathbb{E}\left[\text{Number of times }\hat{A}_{i}\text{ is pulled in round }i\text{ after the first }T_{i}^{\prime}\text{ times}\right]$ (12)
$=$ $\sum_{s=1}^{\infty}\text{Pr}(k\geq s)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\sum_{s=1}^{\infty}\exp\left(-s^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}\leq 1.$
Summing up the three parts of satisficing regret we have that the satisficing regret of round $i$ is still bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Appendix G Proof of Theorem 7
We first prove that the regret incurred in round $i$ is at most
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
The regret incurred when running $\texttt{ALG}(t_{i}^{\prime})$ is bounded by
$\mathbb{E}\left[t_{i}^{\prime}r(A^{*})-\sum_{s=1}^{t_{i}^{\prime}}r(A_{s})\right]\leq C_{1}t_{i}^{\prime\alpha}\log(t_{i}^{\prime})^{\beta}\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
We further have
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Therefore the satisficing regret incurred when pulling $\hat{A}_{i}$ is bounded by
$\left\lceil\dfrac{1}{\gamma_{i}^{2}}\right\rceil\cdot\dfrac{2C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}=\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-1}\log(2/\gamma_{i})^{\beta}\leq\dfrac{3C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since $r(\hat{A}_{i})\leq r(A^{*})<S$, for every $k$, round $i$ is not terminated with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\frac{k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{T_{i}^{\prime}+k}-\sqrt{\frac{k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}^{\prime}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\exp\left(-k^{\zeta}+\log\left(\frac{\zeta}{8\Gamma(\zeta^{-1})}\right)\right)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore expected number of arm pulls of $\hat{A}_{i}$ after pulling $T_{i}^{\prime}$ times is bounded by
$\frac{\zeta}{8\Gamma(\zeta^{-1})}\sum_{s=1}^{\infty}\exp\left(-k^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}\leq 1.$
Summing up the three parts of regret we have that the satisficing regret of round $i$ is bounded by
$\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2/\gamma_{i})^{\beta}.$
Since the length of the time horizon is $T$ time steps, we have
$T\geq\sum_{i}^{i_{\max}}t_{i}^{\prime}\geq\sum_{i}^{i_{\max}}i^{1/\zeta-1}\geq\int_{0}^{i_{\max}}x^{1/\zeta-1}\mathrm{d}x=\zeta i_{\max}^{1/\zeta},$
thus the maximum number of rounds is bounded by $i_{\max}=\lceil(T/\zeta)^{\zeta}\rceil$. Therefore we have that the regret is bounded by
$\mathbb{E}[\texttt{Regret}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\gamma_{i}^{-\alpha/(1-\alpha)}\log(2T)^{\beta}\leq\sum_{i=1}^{i_{\max}}\dfrac{6C_{1}}{(1-\alpha)^{\beta}}i^{\alpha(1-\zeta)/\zeta}\log(2T)^{\beta}$
$\leq$ $\dfrac{6C_{1}}{(1-\alpha)^{\beta}}\log(2T)^{\beta}\int_{1}^{i_{\max}+1}x^{\alpha(1-\zeta)/\zeta}\mathrm{d}x\leq\dfrac{6C_{1}}{(1-\alpha)^{\beta}(\alpha(1-\zeta)/\zeta+1)}(i_{\max}+1)^{\alpha(1-\zeta)/\zeta+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{6C_{1}\zeta^{-\alpha(1-\zeta)-\zeta}\cdot 2^{\alpha(\zeta^{-1}-1)+1}}{(1-\alpha)^{\beta}(\alpha(\zeta^{-1}-1)+1)}\,T^{\alpha(1-\zeta)+\zeta}\log(2T)^{\beta}$
Therefore the regret is bounded by $C_{1}T^{\alpha(1-\zeta)+\zeta}\cdot\textup{polylog}(T)$.
Appendix H Tail Risk Control for Satisficing in Finite-Armed Bandits
In Section 6, we established SELECT-LITE, which is able to attain a constant expected satisficing regret and a light-tailed satisficing regret distribution in the realizable case. However, in the non-realizable case, the standard regret distribution of SELECT-LITE may not be light-tailed. In this section, we focus on the finite-armed bandit setting and establish an algorithm that: 1) attains a constant expected satisficing regret and a light-tailed satisficing regret distribution in the realizable case; 2) attains a sub-linear expected standard regret and a light-tailed standard regret distribution in the non-realizable case.
Similar to Condition 1, we first assume access to a learning oracle that attains a sub-linear expected standard regret and a light-tailed standard regret distribution for finite-armed bandits.
Condition 2.
There exits an finite-armed bandit algorithm $\texttt{ALG}(t)$, functions $\Lambda(K,t),\Lambda^{\prime}(K,t)$ and $\zeta>0$ such that for any instance of $K$-arm bandits and any time horizon $t$,
  1. 1.
    The expected standard regret of $\texttt{ALG}(t)$ is upper bounded by $CK^{\alpha_{1}}t^{\alpha_{2}}\log(t)^{\beta}$, where $C,\alpha_{1},\alpha_{2},\beta$ are constants independent from $K$ and $t$;
  2. 2.
    The probability that the standard regret of $\texttt{ALG}(t)$ exceeds $x$ is upper bounded by $\Lambda(K,t)\exp(-x^{\zeta}/\Lambda^{\prime}(K,t))$;
  3. 3.
    $\Lambda(K,t)$ can be at most polynomially dependent on $t$ and $\Lambda^{\prime}(K,t)$ can be at most poly-logarithmically dependent on $t$.
Without loss of generality we also assume that $\Lambda(K,t)$ and $\Lambda^{\prime}(K,t)$ are monotonically increasing in $t$ for every $K$, and is monotonically decreasing in $x$ for every fixed $t$ and $K$. Condition 2 can be satisfied by existing algorithms, such as the one established in Simchi-Levi et al. (2024).
For finite-armed bandits, we notice that the oracle algorithm is able to output the candidate arm that asymptotically converges to the optimal arm by choosing the one that is the most frequently pulled. Because the arm is already been pulled by at least $t_{i}/K$ times, leveraging this history already yields a tightened confidence interval for it, making forced sampling unnecessary. This holds for both realizable and non-realizable cases.
Therefore, we make three modifications to SELECT to make both satisficing regret distribution and standard regret distribution light-tailed for finite-armed bandits with the help of a light-tailed oracle algorithm. First, we get rid of forced sampling and directly start LCB test after running the oracle. Second, we use the arm that is pulled most frequently during $\texttt{ALG}(t_{i})$ as the candidate arm, and the historical data of this arm being pulled when running $\texttt{ALG}(t_{i})$ can be used in LCB test. Third, we use larger confidence intervals in the LCB test the same as the one in SELECT-LITE. The complete algorithm is provided in Algorithm 3.
Algorithm 3 Light-Tailed Satisficing Regret Minimization for Finite-Armed Bandits (SELECT-LITE+)
Input: time horizon $T$, satisficing level $S$, tail distribution parameter $\zeta$
Set $\gamma_{i}\leftarrow 2^{-i(1-\alpha)/\alpha}$ for all $i=1,2,3\dots$
for round $i=1,2,\dots$ do
     Set $t_{i}\leftarrow\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil$
     Run $\texttt{ALG}(t_{i})$, let $\hat{A}_{i}$ denote the arm that is pulled most frequently along the trajectory during $\texttt{ALG}(t_{i})$.
     Set $\tau_{i}$ be the number of times $\hat{A}_{i}$ is pulled during $\texttt{ALG}(t_{i})$, $k\leftarrow 0$.
     Set $\hat{r}_{i}^{\text{tot}}$ be the sum of observed reward of pulling $\hat{A}_{i}$ for $\tau_{i}$ times during $\texttt{ALG}(t_{i})$.
     while $LCB(\hat{A}_{i})\geq S$ do
         Set $k\leftarrow k+1$
         Pull arm $\hat{A}_{i}$ and observe $Y_{t}$, where $t$ is the current time step
         Update $\hat{r}_{i}^{\text{tot}}\leftarrow\hat{r}_{i}^{\text{tot}}+Y_{t}$
         Set
$LCB(\hat{A}_{i})\leftarrow\dfrac{\hat{r}_{i}^{\text{tot}}}{T_{i}+k}-\sqrt{\frac{2k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{T_{i}+k}}$
     end while
end for
[…]
With the help of a finite-armed bandit learning algorithm with light-tailed standard regret, we use the following theorem to show that SELECT-LITE+ can enjoy a light-tailed satisficing regret distribution in the realizable case and a light-tailed standard regret distribution in the non-realizable case.
Theorem 8.
If $r(A^{*})\geq S$, then under Condition 2, for any $x>0$ the probability that satisficing regret of SELECT-LITE+  exceeds $x$ is bounded by
$\text{Pr}\left(\texttt{Regret}_{S}>x\right)$
$\leq$ $(\alpha_{2}\log_{2}(T)+1)\Lambda(K,T)\exp\left(\dfrac{-x^{\zeta}}{2(\alpha_{2}\log_{2}(T)+1)\Lambda^{\prime}(K,T)}\right)+\exp\left(2i_{0}+4-(x/2)^{\zeta}\right),$
where $i_{0}=O\left(\log_{2}\left(1/\Delta_{S}^{*}\cdot\log(1/\Delta_{S}^{*})^{\beta^{\prime}}\right)\right)$ is a constant independent of $x$.
If $r(A^{*})<S$ and the oracle algorithm ALG is light-tailed, then the probability that standard regret of SELECT-LITE+  exceeds $x$ is bounded by
$\text{Pr}\left(\texttt{Regret}>x\right)$
$\leq$ $(\alpha_{2}\log_{2}(T)+1)\Lambda(K,T)\exp\left(\dfrac{-x^{\zeta}}{2(\alpha_{2}\log_{2}(T)+1)\Lambda^{\prime}(K,T)}\right)+2T^{\alpha_{2}}\exp\left(-(x/2)^{\zeta}\right).$
Proof of Theorem 8 is provided in Section H.1. In the next theorem, we show that similar to SELECT, the satisficing regret of SELECT-LITE+ can be bounded by a constant in the realizable case.
Theorem 9.
If $r(A^{*})\geq S$, then the satisficing regret of SELECT-LITE+ is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\min\left\{C_{1}\left(\dfrac{C_{1}K^{\alpha_{1}+1}}{\Delta_{S}^{*}}\right)^{\frac{\alpha_{2}}{(1-\zeta)(1-\alpha_{2})}}\cdot\textup{polylog}\left(\dfrac{C_{1}K^{\alpha_{1}+1}}{\Delta_{S}^{*}}\right),\,C_{1}T^{\alpha_{2}}\cdot\textup{polylog}(T)\right\}.$
Proof of Theorem 9 is provided in Section H.2. The proof relies on two critical results. In Proposition 7, we show an upper bound on the satisficing regret of round $i$. The proof is provided in Section H.3 of the appendix. In Proposition 8, we also show that once SELECT-LITE+  runs for enough rounds, it is unlikely to start a new round. The proof is provided in Section H.4 of the appendix.
Proposition 7.
If $r(A^{*})\geq S$, then the satisficing regret incurred in round $i$ is bounded by
$\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
Proposition 8.
If $r(A^{*})>S$, then round $i$ terminates within the time horizon with probability at most $1/4$ if $i$ that satisfies
$\dfrac{32C_{1}K^{\alpha_{1}+1}(1-\zeta)^{-\zeta/2}}{(1-\alpha)^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*}\,.$
We also show that in the non-realizable case, SELECT-LITE+ enjoys the same standard regret bound as the oracle shown in Theorem 10. The proof is provided in Section H.5 of the appendix.
Theorem 10.
If $r(A^{*})<S$, the standard regret of SELECT-LITE+is bounded by
$C_{1}K^{\alpha_{1}+1}T^{\alpha_{2}}\cdot\textup{polylog}(T).$
H.1 Proof of Theorem 8
For notational brevity we define $p(x;K,t)=\Lambda(K,t)\exp(-x^{\zeta}/\Lambda^{\prime}(K,t))$. We first bound the tail probability of running the oracle. We define the satisficing regret incurred in the oracle algorithm in each round $i$ as $R_{i}$. Recall that $i_{\max}=\lceil\alpha_{2}\log_{2}(T)\rceil$, then the probability that cumulative satisficing regret incurred in oracle algorithm exceeds $x/2$ is
$\text{Pr}\left(\sum_{i=1}^{i_{\max}}R_{i}>\frac{x}{2}\right)\leq\sum_{i=1}^{i_{\max}}\text{Pr}\left(R_{i}>\frac{x}{2i_{\max}}\right)$ (13)
$\leq$ $\sum_{i=1}^{i_{\max}}p\left(\frac{x}{2i_{\max}};\,K,t_{i}\right)\leq i_{\max}p\left(\frac{x}{2i_{\max}};\,K,T\right)$
$\leq$ $(\alpha_{2}\log_{2}(T)+1)p\left(\frac{x}{2(\alpha_{2}\log_{2}(T)+1)};\,K,T\right)$
The third inequality is because $p(x;\,K,t)$ is monotonically increasing in $t$, and the fourth inequality is because $p(x;\,K,t)$ is monotonically decreasing in $x$.
We then bound the tail probability of LCB test. This is divided into realizable and non-realizable cases. For the realizable case, where Proposition 8 holds, we recall (9), then the probability that satisficing regret incurred in the LCB test in round $i$ exceeds $x/2$ can be bounded by
$\text{Pr}\left(\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/2\right)\leq\text{Pr}\left(k_{i}>x/2\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-2(x/2)^{\zeta}\right).$
Note that $\zeta/\Gamma(\zeta^{-1})\leq 3/2$, $\exp\left(2^{\zeta}\right)\leq e^{2}$. Using Proposition 5, the probability of total regret in the LCB test exceeds $x/2$ can be bounded by
$\text{Pr}\left(\sum_{i=1}^{\infty}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/2\right)=\text{Pr}\left(\sum_{i=1}^{i_{0}+N(1/4)}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/2\right)$ (14)
$\leq$ $\left(1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(2^{\zeta}\right)\right)^{i_{0}}\exp\left(\frac{4}{3}\left(1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(2^{\zeta}\right)\right)-(x/2)^{\zeta}\right)\leq\exp\left(2i_{0}+4-(x/2)^{\zeta}\right)$
For the non-realizable case, the probability that cumulative satisficing regret incurred in oracle algorithm exceeds $x/2$ is upper bounded by
$\text{Pr}\left(\sum_{i=1}^{\infty}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/2\right)\leq\text{Pr}\left(\left(\sum_{i=1}^{i_{\max}}k_{i}\right)^{\zeta}>\left(\frac{x}{2}\right)^{\zeta}\right)$
$\leq$ $\text{Pr}\left(\exp\left(\sum_{i=1}^{i_{\max}}k_{i}^{\zeta}\right)>\exp\left(\left(\frac{x}{2}\right)^{\zeta}\right)\right)\leq\frac{\mathbb{E}\left[\exp\left(\sum_{i=1}^{i_{\max}}k_{i}^{\zeta}\right)\right]}{\exp((x/2)^{\zeta})}\leq\exp\left(-(x/2)^{\zeta}\right)\prod_{i=1}^{i_{\max}}\mathbb{E}\left[\exp\left(k_{i}^{\zeta}\right)\right].$
For each term $\mathbb{E}\left[\exp\left(k_{i}^{\zeta}\right)\right]$, it can be bounded by
$\mathbb{E}\left[\exp\left(k_{i}^{\zeta}\right)\right]=\sum_{s=0}^{\infty}\text{Pr}\left(k_{i}=s\right)\cdot\exp\left(s^{\zeta}\right)\leq 1+\sum_{s=1}^{\infty}\text{Pr}\left(k_{i}=s\right)\cdot\exp\left(s^{\zeta}\right)$
$\leq$ $1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\sum_{s=1}^{\infty}\exp\left(-2s^{\zeta}+s^{\zeta}\right)\leq 1+\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})\leq 2$
which is independent of the round $i$. Then we can bound the probability that cumulative satisficing regret incurred in oracle algorithm exceeds $x/2$ by
$\text{Pr}\left(\sum_{i=1}^{\infty}\max\left\{S-r(\hat{A}_{i}),0\right\}\cdot k_{i}>x/2\right)\leq\exp\left(-(x/2)^{\zeta}\right)\cdot 2^{1+\alpha_{2}\log_{2}(T)}=2T^{\alpha_{2}}\exp\left(-(x/2)^{\zeta}\right).$ (15)
Combining oracle tail probability (13) with the tail probability (14) in the LCB test for the realizable case and the tail probability (15) in the LCB test for the non-realizable case, we have the tail probability to be
$\text{Pr}\left(\texttt{Regret}_{S}>x\right)\leq(\alpha_{2}\log_{2}(T)+1)p\left(\frac{x}{2(\alpha_{2}\log_{2}(T)+1)};\,K,T\right)+\exp\left(2i_{0}+4-(x/2)^{\zeta}\right)$
for the realizable case and tail probability to be
$\text{Pr}\left(\texttt{Regret}>x\right)\leq(\alpha_{2}\log_{2}(T)+1)p\left(\frac{x}{2(\alpha_{2}\log_{2}(T)+1)};\,K,T\right)+2T^{\alpha_{2}}\exp\left(-(x/2)^{\zeta}\right).$
H.2 Proof of Theorem 9
Denote $\beta^{\prime}=\max\{\beta,1/2\}$ and $K_{\alpha,\beta,\zeta}=\dfrac{32C_{1}K^{\alpha_{1}+1}(1-\zeta)^{-\zeta/2}}{(1-\alpha_{2})^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}$, and $i_{0}$ as
$i_{0}=\max\{i:K_{\alpha,\beta,\zeta}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}>\Delta_{S}^{*}\}.$
We first prove an upper bound on $i_{0}$. Denote $\delta_{0}\in(0,1)$ the smallest solution to the following equation
$\delta_{0}^{1-\zeta}\log(2/\delta_{0})^{\beta^{\prime}}=\dfrac{\Delta_{S}^{*}}{K_{\alpha,\beta,\zeta}}.$
Then we have
$\dfrac{1}{\delta_{0}^{1-\zeta}}=\dfrac{K_{\alpha,\beta,\zeta}\log(1/\delta_{0})^{\beta^{\prime}}}{\Delta_{S}^{*}}=\dfrac{(4\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}\log(2/\delta_{0}^{1/4\beta^{\prime}})^{\beta^{\prime}}}{\Delta_{S}^{*}}\leq\dfrac{(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}{\Delta_{S}^{*}(\delta_{0})^{1/4}},$
or equivalently
$\dfrac{1}{\delta_{0}^{1-\zeta}}\leq\left(\dfrac{(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}{\Delta_{S}^{*}}\right)^{4/3},$
therefore
$\dfrac{1}{\delta_{0}^{1-\zeta}}\leq\dfrac{(4(1-\zeta)/3)^{\beta^{\prime}}K_{\alpha,\beta,\zeta}\log({2(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}}/{\Delta_{S}^{*}})^{\beta^{\prime}}}{\Delta_{S}^{*}}$
We write $D_{\alpha,\beta,\zeta}=2(8\beta^{\prime})^{\beta^{\prime}}K_{\alpha,\beta,\zeta}$, $L_{\alpha,\beta,\zeta}=(4(1-\zeta)/3)^{\beta^{\prime}}K_{\alpha,\beta,\zeta}$, then we have
$\delta_{0}^{1-\zeta}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}.$
By definition of $i_{0}$ we also have
$\gamma_{i_{0}}^{1-\zeta}\geq\dfrac{\Delta_{S}^{*}}{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}},$
thus
$i_{0}\leq\dfrac{\alpha_{2}}{(1-\alpha_{2})(1-\zeta)}\log_{2}\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right).$
Using the result of Proposition 7 we have that the satisficing regret by round $i_{0}$ is bounded by
$\sum_{i=1}^{i_{0}}\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\delta_{0})^{\beta}$
$=$ $\sum_{i=1}^{i_{0}}\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i}\log(2/\delta_{0})^{\beta}=\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i_{0}}\log(2/\delta_{0})^{\beta}$
$\leq$ $\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\alpha_{2}}{(1-\alpha_{2})(1-\zeta)}}\log\left(\dfrac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta}.$
By the results of Proposition 8, for every $k\geq 1$, the probability that round $i_{0}+k$ starts within the time horizon is at most $4^{-k+1}$. Therefore the regret incurred after round $i_{0}$ is bounded by
$\sum_{k=1}^{\infty}\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i_{0}+k}\log(2/\gamma_{i})^{\beta}\cdot 4^{-k+1}\leq\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i_{0}}\log(2/\delta_{0})^{\beta}\cdot\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$
$\leq$ $\dfrac{4C_{1}K^{\alpha_{1}+1}M}{(1-\alpha_{2})^{\beta}}\left(\dfrac{L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\frac{\alpha_{2}}{(1-\alpha_{2})(1-\zeta)}}\log\left(\dfrac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta},$
where $M=\sum_{k=1}^{\infty}k^{\beta}\cdot 2^{-k}$ is a constant only dependent on $\alpha,\beta,\zeta$, and the total satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq\dfrac{4C_{1}K^{\alpha_{1}+1}(M+2)}{(1-\alpha_{2})^{\beta}}\left(\dfrac{L_{\alpha,\beta,\zeta}}{\Delta_{S}^{*}}\right)^{\frac{\alpha_{2}}{(1-\alpha_{2})(1-\zeta)}}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\frac{\alpha_{2}\beta^{\prime}}{(1-\alpha_{2})(1-\zeta)}}\log\left(\dfrac{2L_{\alpha,\beta,\zeta}\log(D_{\alpha,\beta,\zeta}/\Delta_{S}^{*})^{\beta^{\prime}}}{\Delta_{S}^{*}}\right)^{\beta}.$
Furthermore, since the length of the time horizon is $T$ time steps, we have $t_{i}\leq T$, thus the maximum number of rounds is bounded by $i_{\max}=\lceil\alpha_{2}\log_{2}(T)\rceil$. By the results of Proposition 7 we have that the satisficing regret is bounded by
$\mathbb{E}[\texttt{Regret}_{S}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2T)^{\beta}\leq\sum_{i=1}^{i_{\max}}\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{16C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}T^{\alpha_{2}}\log(2T)^{\beta}$
H.3 Proof of Proposition 7
Recall that $t_{i}=\lceil\gamma_{i}^{-1/(1-\alpha)}\rceil$, the satisficing regret incurred when running $\texttt{ALG}(t_{i})$ is bounded by
$\mathbb{E}\left[\sum_{s=1}^{t_{i}}\max\{S-r(A_{s}),0\}\right]\leq\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq\dfrac{2C_{1}K^{\alpha_{1}}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
Since $\hat{A}_{i}$ is the most frequently selected arm, it has been pulled for at least $t_{i}/K$ times during $\texttt{ALG}(t_{i})$, and we further have
$\mathbb{E}[\max\{S-r(\hat{A}_{i}),0\}]\leq\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]}{t_{i}/K}\leq\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
If $r(\hat{A}_{i})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{A}_{i}$ for $k$ times (in addition to the number of pulls $\tau_{i}$ during $\texttt{ALG}(t_{i})$) occurs with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{\tau_{i}+k}-\sqrt{\frac{2k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{\tau_{i}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{\tau_{i}+k}-\sqrt{\frac{2k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{\tau_{i}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\exp\left(-2k^{\zeta}+\log\left(\frac{\zeta}{8\Gamma(\zeta^{-1})}\right)\right)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore expected number of arm pulls of $\hat{A}_{i}$ after first pulling it $\tau_{i}$ times is bounded by
$\mathbb{E}\left[\text{Number of times }\hat{A}_{i}\text{ is pulled in round }i\text{ after the first }T_{i}\text{ times}\right]$ (16)
$=$ $\sum_{s=1}^{\infty}\text{Pr}(k\geq s)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\sum_{s=1}^{\infty}\exp\left(-s^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}\leq 1.$
Summing up the two parts of satisficing regret we have that the satisficing regret of round $i$ is bounded by
$\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
H.4 Proof of Proposition 8
Recall from the proof of Proposition 7, we have
$\mathbb{E}[\max\{S-r(\hat{A}_{i}),0\}]\leq\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]}{t_{i}/K}\leq\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
According to the condition of this proposition, we have $i$ satisfies the condition
$\dfrac{32C_{1}K^{\alpha_{1}+1}(1-\zeta)^{-\zeta/2}}{(1-\alpha)^{\beta}}\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*}.$
Note that for any $10<\zeta<$ 1, $(1-\zeta)^{-\zeta/2}>1$ and $\gamma_{i}^{1-\zeta}>\gamma_{i}$. Therefore $i$ satisfies
$\dfrac{32C_{1}K^{\alpha_{1}+1}}{(1-\alpha)^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\max\{\beta,1/2\}}\leq\Delta_{S}^{*},$
by the Markov’s inequality we have
$\text{Pr}(r(A^{*})-r(\hat{A}_{i})\geq\Delta_{S}^{*}/2)\leq\text{Pr}\left(r(A^{*})-r(\hat{A}_{i})\geq\dfrac{16C_{1}K^{\alpha_{1}+1}}{(1-\alpha)^{\beta}}\log(2/\gamma_{i})^{\beta}\right)\leq\dfrac{1}{8}.$ (17)
Now, if $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$, which happens with probability at least 7/8 according to the above inequality, then $r(\hat{A}_{i})\geq S+\Delta_{S}^{*}/2$ holds. On the other hand, if $i$ satisfies the condition stated in the proposition, then
$(1-\zeta)^{-\zeta/2}\cdot\max\left\{2,\sqrt{\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}\right\}\gamma_{i}^{1-\zeta}\log(1/\gamma_{i})^{1/2}\leq\Delta_{S}^{*}/16.$
Note that for $f(y)=y^{\zeta}/(\tau_{i}+y)$, it reaches its maximum when
$\frac{\mathrm{d}}{\mathrm{d}y}f(y)=\frac{\zeta y^{\zeta-1}\tau_{i}+y^{\zeta}(\zeta-1)}{(\tau_{i}+y)^{2}}=0.$
Therefore the maximum of $f(y)$ is reached at $\zeta\tau_{i}/(1-\zeta)$. For simplicity, we set $D=\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)$. The confidence radius for any $k$ is bounded by
$\sqrt{\frac{2k^{\zeta}+D}{\tau_{i}+k}}\leq\sqrt{\frac{\max\{D,3\}k^{\zeta}}{\tau_{i}+k}}\leq\sqrt{\frac{\max\{D,2\}\left(\frac{\zeta\tau_{i}}{1-\zeta}\right)^{\zeta}}{\tau_{i}+\left(\frac{\zeta\tau_{i}}{1-\zeta}\right)}}$
$\leq$ $\sqrt{\frac{\max\{D,2\}(1-\zeta)^{-\zeta}\cdot\zeta^{\zeta}\tau_{i}^{\zeta}}{\tau_{i}}}\leq\max\left\{\sqrt{D},2\right\}(1-\zeta)^{-\zeta/2}\gamma_{i}^{1-\zeta}\leq\frac{\Delta_{S}^{*}}{16}.$
Suppose round $i$ is terminated for some $k$, then
$\dfrac{\hat{r}_{i}^{\text{tot}}}{\tau_{i}+k}-\sqrt{\frac{2k^{\zeta}+D}{\tau_{i}+k}}<S=r(A^{*})-\Delta_{S}^{*}\leq r(\hat{A}_{i})-\dfrac{\Delta_{S}^{*}}{2}\leq r(\hat{A}_{i})-2\sqrt{\frac{2k^{\zeta}+D}{\tau_{i}+k}},$
or equivalently
$\dfrac{\hat{r}_{i}^{\text{tot}}}{\tau_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\frac{2k^{\zeta}+D}{\tau_{i}+k}}.$
By the Hoeffding’s inequality we have
$\text{Pr}\left(\dfrac{\hat{r}_{i}^{\text{tot}}}{\tau_{i}+k}\leq r(\hat{A}_{i})-\sqrt{\frac{2k^{\zeta}+D}{\tau_{i}+k}}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-2k^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore conditioned on $r(\hat{A}_{i})\geq r(A^{*})-\Delta_{S}^{*}/2$ the probability that round $i$ is terminated by the end of the horizon is bounded by
$\sum_{k=1}^{\infty}\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$ (18)
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}.$
Combining (17) and (18) we conclude that the probability that round $i$ terminates by the end of the time horizon is bounded by $1/4$.
H.5 Proof of Theorem 10
We first prove that the regret incurred in round $i$ is at most
$\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
The standard regret incurred when running $\texttt{ALG}(t_{i})$ is bounded by
$\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]\leq\dfrac{2C_{1}K^{\alpha_{1}}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
Since $\hat{A}_{i}$ is the most frequently selected arm, it has been pulled for at least $t_{i}/K$ times during $\texttt{ALG}(t_{i})$, and we further have
$\mathbb{E}[r(A^{*})-r(\hat{A}_{i})]\leq\dfrac{\mathbb{E}\left[t_{i}r(A^{*})-\sum_{s=1}^{t_{i}}r(A_{s})\right]}{t_{i}/K}\leq\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}\log(2/\gamma_{i})^{\beta}.$
Since $r(\hat{A}_{i})\leq r(A^{*})<S$, then for every $k$, round $i$ is not terminated after pulling arm $\hat{A}_{i}$ for $k$ times (in addition to the number of pulls $\tau_{i}$ during $\texttt{ALG}(t_{i})$) occurs with probability at most
$\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{\tau_{i}+k}-\sqrt{\frac{2k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{\tau_{i}+k}}>S\right)$
$\leq$ $\text{Pr}\left(\dfrac{\hat{r}^{\text{tot}}_{i}}{\tau_{i}+k}-\sqrt{\frac{2k^{\zeta}+\log\left(8\zeta^{-1}\Gamma(\zeta^{-1})\right)}{\tau_{i}+k}}>r(\hat{A}_{i})\right)$
$\leq$ $\exp\left(-2k^{\zeta}+\log\left(\frac{\zeta}{8\Gamma(\zeta^{-1})}\right)\right)=\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right).$
Therefore expected number of arm pulls of $\hat{A}_{i}$ after first pulling it $\tau_{i}$ times is bounded by
$\sum_{s=1}^{\infty}\frac{\zeta}{8\Gamma(\zeta^{-1})}\exp\left(-k^{\zeta}\right)\leq\frac{\zeta}{8\Gamma(\zeta^{-1})}\int_{0}^{\infty}\exp\left(-x^{\zeta}\right)\mathrm{d}x$
$\leq$ $\frac{\zeta}{8\Gamma(\zeta^{-1})}\cdot\zeta^{-1}\Gamma(\zeta^{-1})=\frac{1}{8}\leq 1.$
Summing up the two parts of standard regret we have that the standard regret of round $i$ is bounded by
$\dfrac{2C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2/\gamma_{i})^{\beta}.$
Since the length of the time horizon is $T$ time steps, we have $t_{i}\leq T$, thus the maximum number of rounds is bounded by $i_{\max}=\lceil\alpha_{2}\log_{2}(T)\rceil$. Therefore, we have that the standard regret is bounded by
$\mathbb{E}[\texttt{Regret}]\leq$ $\sum_{i=1}^{i_{\max}}\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}\gamma_{i}^{-\alpha_{2}/(1-\alpha_{2})}\log(2T)^{\beta}\leq\sum_{i=1}^{i_{\max}}\dfrac{4C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}2^{i+1}\log(2T)^{\beta}$
$\leq$ $\dfrac{16C_{1}K^{\alpha_{1}+1}}{(1-\alpha_{2})^{\beta}}T^{\alpha_{2}}\log(2T)^{\beta}$
Therefore the regret is bounded by $C_{1}K^{\alpha_{1}+1}T^{\alpha_{2}}\cdot\text{polylog}(T)$.
Appendix I Choices of $C_{1},\alpha,\beta$ for Several Bandit Optimization Problem Classes
In this section, we provide a detailed discussion on the choice of $C_{1},\alpha,\beta$ for several bandit optimization problem classes. In three cases
I.1 Finite-armed Bandits
Consider any instance of finite-armed bandits with mean reward bounded by $[0,1]$. By Theorem 7.2 of Lattimore and Szepesvári (2020), for any given time horizon $t\geq 2$, the standard regret of the UCB algorithm is upper bounded by
$8\sqrt{Kt\log(t)}+3\sum_{k=1}^{K}(r(k^{*})-r(k))\leq 8\sqrt{Kt\log(t)}+3K,$
where $K\geq 2$ is the number of arms and $k^{*}$ is the optimal arm.
When $t>K$, the regret is upper bounded by $8\sqrt{Kt\log(t)}+3K\leq 11\sqrt{Kt\log(t)}$, where the last inequality is due to $t>K\geq 2$, thus $t\geq 3$ and $\log(t)\geq\log(3)>1$; When $t\leq K$, the regret is upper bounded by $t\leq\sqrt{Kt}\leq 11\sqrt{Kt\log(2)}\leq 11\sqrt{Kt\log(t)}$.
Thus for any given time horizon $t\geq 2$, the regret of the UCB algorithm is bounded by $11\sqrt{Kt\log(t)}$. Therefore, by setting $C_{1}=11\sqrt{K}$ and $\alpha=\beta=1/2$ and using UCB algorithm as the oracle, Condition 1 is satisfied for finite-armed bandits.
I.2 Concave Bandits
Consider any concave bandits with $\mathcal{A}\subseteq\mathbb{R}^{d}$ being a compact and convex set contained in a Euclidean ball with radius $1$, and reward function being concave and $1$-Lipschitz. By Theorem 2 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 2 in Agarwal et al. (2011) is upper bounded by
$768d^{3}\sqrt{t}\log(t)^{2}\left(\dfrac{2d^{2}\log(d)}{c_{2}^{2}}+1\right)\left(\dfrac{4d^{7}c_{1}}{c_{2}^{3}}+\dfrac{d(d+1)}{c_{2}}\right)\left(\dfrac{4c_{1}d^{4}}{c_{2}^{2}}+11\right),$
where $c_{1},c_{2}$ are algorithm hyperparameters that satisfies $c_{1}\geq 64$, $c_{2}\leq 1/32$. Therefore Condition 1 is satisfied for concave bandits by using Algorithm 2 in Agarwal et al. (2011) and setting $\alpha=1/2$, $\beta=2$ and
$C_{1}=768d^{3}\left(\dfrac{2d^{2}\log(d)}{c_{2}^{2}}+1\right)\left(\dfrac{4d^{7}c_{1}}{c_{2}^{3}}+\dfrac{d(d+1)}{c_{2}}\right)\left(\dfrac{4c_{1}d^{4}}{c_{2}^{2}}+11\right),$
Furthermore, if the instance of concave bandit is one-dimensional, i.e., $\mathcal{A}=[0,1]$, by Theorem 1 of Agarwal et al. (2011), for any given time horizon $t\geq 2$, the standard regret of Algorithm 1 in Agarwal et al. (2011) is bounded by
$108\sqrt{t\log(t)}\cdot\log_{4/3}(t)=108/\log(4/3)\sqrt{t}\cdot\log(t)^{3/2}.$
Therefore Condition 1 is also satisfied for one-dimensional concave bandits by using Algorithm 1 in Agarwal et al. (2011) and setting $C_{1}=108/\log(4/3)$, $\alpha=1/2$, $\beta=3/2$.
I.3 Lipschitz Bandits
Consider any Lipschitz bandits with arm set being $[0,1]^{d}$, reward function being $L$-Lipschitz and mean reward bounded on $[0,1]$ (here we define Lipschitz continuity in the sense of $\infty$-norm). By Section 2.2 of Bubeck et al. (2011), if the regret of the UCB algorithm for finite-armed bandits with $K$ arms is bounded by $c_{\text{ALG}}\sqrt{Kt\log(t)}$ for any $t\geq 2$, then for any given time horizon $t\geq 2$, the standard regret of the uniformly discretized UCB algorithm introduced by Bubeck et al. (2011) (hereafter referred to as the “Uniform UCB”) is upper bounded by
$(1+c_{\text{ALG}})L^{d/(d+2)}t^{(d+1)/(d+2)}\cdot\sqrt{\log(t)}.$
Using the results in Section I.1, we have $c_{\text{ALG}}=11$, thus the regret of the Uniform UCB algorithm is bounded by
$12L^{d/(d+2)}t^{(d+1)/(d+2)}\cdot\sqrt{\log(t)}.$
Therefore, by setting $C_{1}=12L^{d/(d+2)}$, $\alpha=(d+1)/(d+2)$, $\beta=1/2$ and using Uniform UCB as the oracle, Condition 1 is satisfied for Lipschitz bandits.
Appendix J Ablation Study for SELECT
In this section, we use an ablation study to demonstrate that all three steps of SELECT play a vital role in obtaining a constant satisficing regret. Specifically, we compare the performance of SELECT with the following adapted versions of SELECT:
  • SELECT without Step 1: instead of identifying candidate satisficing arms using a learning oracle, the candidate satisficing arm is drawn uniformly at random from the arm set before we proceed to forced sampling and LCB test in each round;
  • SELECT without Step 2: After identifying a candidate satisficing arm from Step 1 in each round, the forced sampling in Step 2 is skipped and a LCB test on the candidate satisficing arm is immediately started;
  • SELECT without Step 3: Each round is terminated after running Step 1 and Step 2 without entering the LCB test in Step 3.
Setup: We conduct the ablation test using an instance of the Lipschitz bandit. We use the reward function defined in Section 7.3, and set the satisficing threshold $S=0.7$. The learning oracle we use in Step 1 is Uniform UCB introduced in Bubeck et al. (2011).
Results: The results are provided in Figure 10. One can see that after removing any of the three steps, our algorithm is unable to maintain constant satisficing regret in the realizable case. Specifically, Step 1 serves the purpose of efficiently identifying a candidate satisficing arm. If Step 1 is replaced by randomly drawing candidate satisficing arms from the arm set, then the satisficing regret is almost linear in $T$, indicating that the algorithm struggles to identify satisficing arms without the help of the learning oracle. Step 2 and Step 3 combined mainly serve the purpose of exploiting the candidate satisficing arm and at the same time, testing whether it is truly satisficing. Without either Step 2 or Step 3, the satisficing regret starts a round of rapid increase whenever the time horizon $T$ exceeds certain thresholds, thus fail to maintain constant satisficing regret. This indicates that without either Step 2 or Step 3, the algorithm almost always terminates a round shortly after completing Step 1 even if the candidate satisficing arm obtained from Step 1 is indeed satisficing. We refer to Remark 3 for a detailed discussion on the reason for this. Therefore we conclude that all three steps of SELECT play a vital role in obtaining a constant satisficing regret bound in the realizable case.
Figure 10: Results of the ablation test
Appendix K Robustness on the choice of $\gamma_{i}$
(a) Realizable case
(b) Non-realizable case
Figure 11: Performance of SELECT with different choice of $\gamma_{i}$
In Algorithm 1, we define $\gamma_{i}=2^{-i(1-\alpha)/\alpha}$ for all $i\in\mathbb{Z}_{+}$, but all our results still hold if all $\gamma_{i}$ are multiplied by a constant. In this section, we conduct numerical experiments on testing the robustness of SELECT in the choice $\gamma_{i}$.
Setup: We conduct the experiment on the same instance of Lipschitz bandit in Section 7.3. We multiply the $\{\gamma_{i}\}$ in Algorithm 1 with a constant $\lambda$, and vary $\lambda=0.5,1,2,5$.
Results: The result of the realizable case is provided in Figure 11(a), and the result of the non-realizable case is provided in Figure 11(b). One can see from the results that different choices of $\gamma_{i}$ have limited impact on the empirical performance of SELECT, both in the realizable and non-realizable case. Therefore the empirical performance of SELECT is robust in the choice of $\gamma_{i}$.